### v1

# ClinicalReTrial: A Self-Evolving AI Agent for Clinical Trial Protocol Optimization

Abstract
Clinical trial failure remains a central bottleneck in drug development, where minor protocol design flaws can irreversibly compromise outcomes despite promising therapeutics. Although cutting-edge AI methods achieve strong performance in predicting trial success, they are inherently reactive for merely diagnosing risk without offering actionable remedies once failure is anticipated. To fill this gap, this paper proposes ClinicalReTrial, a self-evolving AI agent framework that addresses this gap by casting clinical trial reasoning as an iterative protocol redesign problem. Our method integrates failure diagnosis, safety-aware modification, and candidate evaluation in a closed-loop, reward-driven optimization framework. Serving the outcome prediction model as a simulation environment, ClinicalReTrial  enables low-cost evaluation of protocol modifications and provides dense reward signals for continuous self-improvement. To support efficient exploration, the framework maintains hierarchical memory that captures iteration-level feedback within trials and distills transferable redesign patterns across trials. Empirically, ClinicalReTrial  improves $83.3\%$ of trial protocols with a mean success probability gain of $5.7\%$, and retrospective case studies demonstrate strong alignment between the discovered redesign strategies and real-world clinical trial modifications.
ClinicalReTrial: A Self-Evolving AI Agent for Clinical Trial Protocol Optimization
Sixue Xing1, Xuanye Xia2, Kerui Wu3, Meng Jiang1, Jintai Chen4, Tianfan Fu5 1University of Notre Dame, South Bend, IN, USA 2Georgia Institute of Technology, Atlanta, GA, USA 3University of Massachusetts Amherst, Amherst, MA, USA 4Hong Kong University of Science and Technology (Guangzhou), Guangzhou, Guangdong, China 5Nanjing University, Nanjing, Jiangsu, China Correspondence: futianfan@gmail.com
1 Introduction
Clinical trials represent the most critical and expensive phase in drug discovery, with an estimated cost of \$2.6 billion (dimasi2016innovation) per approved drug, and low success rates of approximately 10-20% (yamaguchi2021approval). Serving as documented plan that specifies the study’s objectives, clinical trial protocols involve complex, interdependent design choices (getz2017trends), such as eligibility criteria, dosing strategies, and endpoint definitions, where small design flaws can propagate into irreversible failure. These challenges motivate the use of AI systems zhang2023harnessing that can reason over high-dimensional trial designs, leverage historical evidence, and systematically assess failure risks at scale.
Recent advances in AI have enabled increasingly accurate prediction of clinical trial outcomes. For example, lo2019machine uses structured metadata to model success likelihood; fu2022hint; chen2024uncertainty; chen2025trialbench integrate heterogeneous data sources using architectures including graph neural networks and hierarchical attention mechanisms to achieve strong predictive performance; yue2024clinicalagent; liu2025autoct incorporate Large Language Models (LLMs) and external knowledge bases to enhance reasoning and explainability in trial outcome prediction.
Despite their success, existing approaches are inherently reactive in nature: they operate on a fixed clinical trial protocol and produce a prediction or post-hoc explanation of trial success or failure. However, these methods do not address a more practically consequential problem: they are unable to respond to a determined trial failure due to the lack of actionable interventions. In real-world drug discovery, stakeholders require not only assessments of failure risk, but also actionable guidance on protocol redesign, including principled modifications or augmentations informed by the identified sources of risk.
To bridge this gap, in this work, we propose ClinicalReTrial, a self-evolving AI agent framework that moves beyond static prediction toward actionable intervention via end-to-end trial protocol optimization, while continuously improves its redesign policies. The framework instantiates a coordinated multi-agent pipeline that performs failure diagnosis, protocol redesign, and candidate evaluation, with domain knowledge and safety awareness embedded at each decision stage. Beyond a single optimization run, we adopt the prediction model as a simulation environment to provide rewards for continuous self-improvement. Specifically, ClinicalReTrial  maintains local memory to accumulate iteration-level feedback and reward-attributed modification outcomes for within-trial adaptation, while a global memory distills transferable redesign patterns across trials to enable warm start initialization and exploration calibration. Through this hierarchical learning structure and reward-driven closed-loop optimization, ClinicalReTrial  systematically explores the protocol modification space and learns to identify high-impact interventions that improve clinical trial success probability.
Experimentally, our prediction model demonstrated the strongest performance (PR-AUC >0.75>0.75), allowing it to serve as a reliable simulation environment for evaluation and agent optimization. In the trial redesign experiments, ClinicalReTrial successfully improve $89.3\%$ of trial protocols with mean probability gain $\Delta p=5.7\%$, achieved at negligible cost ($\mathdollar 0.12/trial$). We further conduct multiple real-world retrospective case studies. Impressively, the redesigns generated by ClinicalReTrial  exhibit strong strategic alignment with independently derived real-world trial modifications, highlighting the potential of self-evolving AI agents to support principled, clinically grounded trial redesign.
Main contributions are listed as follows: (1) (to the best of our knowledge) We are the first to formulate clinical trial optimization as an AI-solvable and in silico-verifiable problem. (2) We propose a multi-agent pipeline with domain knowledge that decomposes clinical trial protocol optimization into analysis, augmentation, and evaluation. (3) We develop a simulation-driven clinical trial optimization framework with hierarchical memory utilization for continuous self-improvement.
2 Related Work
Early efforts employed classical machine learning (logistic regression (LR), random forests) on expert-curated features (gayvert2016clinical; lo2019machine), establishing feasibility but lacking multi-modal data integration. Deep learning approaches addressed this: fu2022hint proposed HINT, integrating drug molecules, ICD-10 codes, and eligibility criteria; chen2024uncertainty added uncertainty quantification; wang2024twin designed LLM-based patient-level digital twins; chen2025trialbench released a standardized TrialBench with multi-modal baselines. while maintaining competitive performance. Recent LLM approaches demonstrate medical reasoning (singhal2023medpalm), enhanced via retrieval-augmented generation (lewis2020rag) with databases like DrugBank (wishart2018drugbank), Hetionet (himmelstein2017hetionet), and domain-adapted encoders like BioBERT (lee2020biobert). Building on this, yue2024clinicalagent introduced ClinicalAgent, decomposing prediction into specialized sub-task agents with ReAct reasoning (yao2023react). liu2025autoct proposed AutoCT for autonomous feature engineering via Monte Carlo Tree Search (chi2024sela).
However, these methods function as discriminators mapping protocols to success probabilities without explaining why failures occur or how to modify protocols. First to formulate generative optimization, our multi-agent architecture leverages chain-of-thought (wei2022chain), and least-to-most prompting (zhou2023least) for hierarchical problem decomposition.
3 Methodology
Overview
We propose ClinicalReTrial, a self-improving multi-agent system that redesigns failed clinical trials through reward-driven iterative optimization. First, Section 3.1 formulates the clinical trial optimization problem. Then, Section 3.2 details the agent components, Section 3.3 presents the hierarchical learning mechanisms, and Section 3.4 describes the knowledge retrieval system. For ease of exposition, Figure 1 illustrates the whole process and Algorithm 1 formalizes the iterative optimization procedure.
Figure 1: ClinicalReTrial Agent architecture. The system operates through iterative refinement: agents analyze failures, generate augmentations, and receive rewards from the simulation environment. Modifications are distilled into structured knowledge that guides subsequent iterations, enabling progressive improvement.
Algorithm 1 Clinical trial protocol optimization with hierarchical learning.
0: Failed trial $T_{0}$, failure mode $y$, global memory $\mathcal{M}^{\text{global}}$
0: Optimized protocol $T^{*}$, best reward $r_{\text{best}}$
1: Initialize: $r_{\text{best}}\leftarrow 0$, $T^{*}\leftarrow T_{0}$, $\mathcal{M}_{0}^{\text{local}}\leftarrow\emptyset$
2:for $t=1$ to $N_{\text{max}}$ do
[…]
6:  if $r_{\text{max}}>r_{\text{best}}$ then $r_{\text{best}}\leftarrow r_{\text{max}}$, $T^{*}\leftarrow\arg\max_{T^{\prime}\in\text{explored}}r(T^{\prime})$
[…]
8:end for
[…]
10:return $T^{*}$, $r_{\text{best}}$
3.1 Problem: Clinical Trial Optimization
Let $T_{0}=\{e_{1},e_{2},\ldots,e_{K}\}$ denote a clinical trial protocol decomposed into $K$ modifiable elements (e.g., eligibility criteria, dosage regimens, endpoint definitions), where a prediction model $f_{\theta}:\mathcal{T}\rightarrow[0,1]$ assigns success probability $p_{0}=f_{\theta}(T_{0})$. Each element $e_{i}$ admits a set of augmentations $\mathcal{A}_{i}=\{a_{i1},a_{i2},\ldots,a_{im_{i}}\}$ representing clinically valid modifications, constructing candidate protocols $T^{\prime}=T_{0}\oplus\mathcal{S}$ by recombining augmented elements with the original protocol. Given the exploration set of redesigned protocol candidates $\mathbb{T}=\{T^{\prime}_{1},T^{\prime}_{2},\ldots,T^{\prime}_{N}\}$ evaluated through the model to obtain predicted probabilities, each $p^{\prime}_{j}=f_{\theta}(T^{\prime}_{j})$, the optimization objective seeks for maximizing success probability ($p^{*}$), with overall improvement measured by $\Delta p=p^{*}-p_{0}$ where $p^{*}=f_{\theta}(T^{*})$.
3.2 Agent Pipeline
3.2.1 Analysis: Structured Failure Diagnosis
The Analysis Agent identifies modification targets within failed trial protocols. Given a failed protocol and prior failure modes: {Poor Enrollment, Safety/Adverse Effect, Drug Lack of Efficacy}. the agent produces a prioritized set of modification, targeting the protocol feature under consideration. Then specifies the action strategy, {Delete, Modify, Add} with confidence score.
Protocol Taxonomy.
Protocol features are classified by catagories: eligibility criteria into participation barriers, safety exclusions, selection criteria, and enrichment criteria; dosage/outcomes by safety risk, failure contribution, and modification efficacy. This taxonomy guides action selection given the observed failure reason and according analysis.
Action Determination.
The agent extracts failure signatures via action category alignment and confidence scoring, pre-calibrated using historical modification success patterns. At iteration $t=1$, the agent receives warm start guidance from cross-trial memory (§3.3.3); from $t\geq 2$, it incorporates performance patterns from prior iterations (§3.3.2). These insights yield prioritized modification targets balancing domain knowledge with empirical feedback, which guide the Augmentation Agent (§3.2.2) to generate concrete modifications.
3.2.2 Augmentation: Structurally Adaptive Variant Generation
The Augmentation Agent translates diagnostic insights from the Analysis Agent into diverse design refinements that address identified weaknesses while preserving clinical validity.
Action-specific Variant Generation.
The agent employs action-specific logic: Delete critical failure factors while preserving safety; Modify adjusts thresholds or operationalizes vague terms; Add introduces biomarker enrichment or contraindication criteria. Multiple variants per target enable exploration, with candidates proceeding to validation (§3.2.3).
3.2.3 Self-validation: Progressive Safety Assurance
To make sure ClinicalReTrialAgent’s proposed augmentations satisfy clinical safety standards, the system employs LLM-as-a-Judge (zheng2023judgingllmasajudgemtbenchchatbot) through hierarchical validation stages.
Autonomous Safety Validation.
The agent checks and prunes unsafe modifications (dosage changes, population shifts, contraindications) and autonomously retrieves evidence from DrugBank, Disease Database, and PubMed (§3.4) when parametric knowledge is insufficient. Validated candidates that passed all stages of progressive filtering proceed to the Exploration Orchestrator (§3.2.4) for simulation-based evaluation and reward assignment.
3.2.4 Exploration Orchestrator with Simulation Environment
The Exploration Orchestrator combines validated augmentations with the original trial protocol into complete redesigned trial candidates, each evaluated through simulation based assessment with outcome probability assigned that guides augmentation rewards and hierarchical learning (§3.3).
Search Strategy.
Candidate trials are formed by combining validated augmentations across orginal protocols, where Beam search is used to reduce exponential complexity to approximately quadratic while maintaining candidate quality.
Simulation Environment.
To provide reliable feedback for agent generated augmentations, we train model that predicts trial candidates’ outcome probabilities from encoded trial features, serving as simulation environment that enables rapid evaluation for thousands of redesigns without conducting actual clinical trials, guiding the agent system toward promising protocol optimization. The improvement in predicted success probability, from the original trial to the redesigned candidate, serves as the reward signal.
3.3 Hierarchical Learning System
Our framework enables progressive improvement through hierarchical knowledge consolidation operating at two temporal scales: within-trial learning accumulates local memory $\mathcal{M}_{t}^{\text{local}}$ across iterations for trial-specific refinement, while cross-trial learning maintains global memory $\mathcal{M}^{\text{global}}$ to transfer successful patterns across the trial corpus.
3.3.1 Redesign Reward
To identify which individual modifications drive improvement, we decompose protocol-level outcome probabilities into augmentation-level rewards. The Exploration Orchestrator first evaluates combined trial variants via prediction, then attributes credit to individual modifications. For each augmentation $m$, we compute its marginal contribution across the explored combinatorial space:
$r(m)=\mathbb{E}_{T^{\prime}\ni m}[p(T^{\prime})]-\mathbb{E}_{T^{\prime}\not\ni m}[p(T^{\prime})].$ (1)
The complete reward distribution $\mathcal{R}_{t}=\{(m_{i},r(m_{i}),v_{i})\}$ encompasses all augmentations with their rewards and validation status, enabling performance-stratified knowledge extraction from both successful modifications and contraindicated patterns.
3.3.2 Within-trial Learning: Iterative Optimization
Knowledge distillation operates on the complete reward distribution $\mathcal{R}_{t}$, serving as short-term memory. Strategic-level knowledge $\mathcal{K}^{s}_{t}$ guides which modification categories merit prioritization. Tactical modification examples $\mathcal{K}^{t}_{t}$ are used to calibrate how modifications should be formulated.
Agent Integration.
The Analysis Agent partitions aspects into three coverage sets: previously failed modifications (zero confidence), previously successful modifications (diversification penalty to avoid over-exploitation), and unexplored modifications (exploration bonus). Confidence scores combine coverage-based adjustments with action-type success rates from historical trials. The Augmentation Agent samples performance-stratified exemplars for few-shot prompting, while the Exploration Agent maintains a redesign pool of top-quartile modifications ($r>0$, 75th percentile) for combinatorial search reuse.
3.3.3 Cross-trial Learning: Global Memory
Global memory $\mathcal{M}^{\text{global}}$ maintains generalizable patterns through two representations: Qualitative strategic guidance (aspect-level recommendations extracted via LLM synthesis from high performing redesign patterns) provides warm start initialization for the Analysis Agent at iteration $t=1$, where iteration feedback is absence; Quantitative statistical signatures (mean reward, variance, modification success rate) enable the Augmentation Agent to continuously calibrate exploration intensity, scaling generation count inversely with historical success rates and proportionally to pattern variance. After each trial converges, patterns distilled from the final iteration, enabling systematic knowledge transfer where each trial benefits from and contributes to the evolving global memory.
3.4 Database Retrieval with Self-reflective Validation
The multi-agent system augments embedded parametric knowledge with targeted retrieval from curated biomedical databases: DrugBank (wishart2018drugbank) with pharmacological profiles including toxicity, metabolism, contraindications; Disease Database (chen2024cod) that contains diagnostic criteria, symptomatology, risk factors; and PubMed Abstract spanning 1975-2025. Retrieval employs dense embeddings (BioBERT (lee2020biobert) for drugs/diseases, PubMedBERT (gu2021domain) for literature) with FAISS indexing. Retrieved results undergo LLM-driven validation, filtering tangential content, while enforcing strict temporal constraints that limit PubMed queries to prevent outcome leakage.
4 Experiment
We evaluate ClinicalReTrial across two dimensions: (1) simulation environment performance, validating that GBDT-based outcome predictors achieve sufficient accuracy to serve as reliable feedback oracles, and (2) ClinicalReTrialAgent optimization quality, demonstrating that our multi-agent system successfully redesigns failed trials through iterative learning.
4.1 Experimental Setup
Our system is built on GPT-4o-mini and evaluated on failed clinical trials from the TrialBench dataset (chen2025trialbench). Using 20769 annotated Phase I-IV trials, we encode multi-modal features into 6,173-dimensional embeddings (details in Appendix A.1.2) and train LightGBM (ke2017lightgbm) classifiers to predict trial outcome $\hat{y}\in[0,1]$. We follow TrialBench’s train-test split, further splitting the training set 8:2 for training-validation. Due to computational constraints, we evaluate the agent on a stratified sample of 60 failed trials from the test set (20 enrollment, 20 safety, 20 efficacy failures) representing diverse trial phases (Data detail in Appendix A.1.1). The agent operates with a 5-iteration budget. We employ an adaptive exploration strategy: when the estimated factorial space <1,000<1,000 combinations, we perform exhaustive exploration to find the optimal solution; otherwise use beam search with width $k=8$. We measure effectiveness through predicted probability improvement, threshold achievement rate, and convergence efficiency.
4.2 Simulation Environment Performance
Our Simulation Environment’s model (GBDT) is compared against Baseline approaches (TrialBench (chen2025trialbench) and HINT (fu2022hint), prior SOTA systems operating on original TrialBench features), and Logistic Regression also trained on the same encodes (Appendix A.1.2).
Failure-specific Prediction.
Implemented in ClinicalReTrialAgent, the simulation environment must correctly predict specific failure outcomes (enrollment, safety, efficacy). We train three independent binary GBDT classifiers on our encoded features, each targeting one failure detection task against success. Table 1, 2 and 3 report comprehensive metrics across our models and baseline approaches re-trained for the binary task: Poor Enrollment, Safety/Adverse Effect and Lack of Efficacy prediction. Our model achieves PR-AUC >0.75>0.75 across all failure modes, meeting the threshold for reliable discriminative feedback. All models achieve Failure Detection Rates of 70-74% at task-specific thresholds ($p\geq 0.6$ for enrollment, $p\geq 0.9$ for safety, $p\geq 0.85$ for efficacy). This ensure that predicted probability shifts $\Delta p>0.03$ reliably indicate improved trial designs (Appendix A.1.1).
Table 1: Performance on Poor Enrollment prediction.
Model ROC-AUC PR-AUC Fail Det.
TrialBench 0.613 $\pm$ 0.007 0.626 $\pm$ 0.011 0.525 $\pm$ 0.013
HINT 0.534 $\pm$ 0.010 0.613 $\pm$ 0.012 0.580 $\pm$ 0.018
Logistic Reg. 0.622 $\pm$ 0.010 0.696 $\pm$ 0.012 0.669 $\pm$ 0.012
GBDT (ours) 0.676 $\pm$ 0.009 0.754 $\pm$ 0.010 0.740 $\pm$ 0.012
Table 2: Performance on Drug Adverse Effect prediction.
Model ROC-AUC PR-AUC Fail Det.
TrialBench 0.587 $\pm$ 0.017 0.892 $\pm$ 0.006 0.427 $\pm$ 0.035
HINT 0.513 $\pm$ 0.014 0.882 $\pm$ 0.009 0.459 $\pm$ 0.031
Logistic Reg. 0.612 $\pm$ 0.018 0.909 $\pm$ 0.008 0.422 $\pm$ 0.029
GBDT (ours) 0.656 $\pm$ 0.018 0.925 $\pm$ 0.007 0.695 $\pm$ 0.028
Table 3: Performance on Drug Efficacy prediction.
Model ROC-AUC PR-AUC Fail Det.
TrialBench 0.692 $\pm$ 0.012 0.862 $\pm$ 0.006 0.565 $\pm$ 0.020
HINT 0.559 $\pm$ 0.013 0.841 $\pm$ 0.008 0.525 $\pm$ 0.021
Logistic Reg. 0.665 $\pm$ 0.015 0.886 $\pm$ 0.009 0.549 $\pm$ 0.025
GBDT (ours) 0.746 $\pm$ 0.013 0.914 $\pm$ 0.007 0.725 $\pm$ 0.021
TrialBench Benchmark.
We further validate the same model architecture against existing benchmarks on the TrialBench 4-class classification task (predicting Success, Enrollment Failure, Safety Failure, or Efficacy Failure). Table 4 shows the Simulation Environment’s base model in our simulation environment performed best over all baselines, with higher ROC-AUC of 0.06 to 0.19.
Category Model Phase 1 Phase 2 Phase 3 Phase 4
ROC-AUC PR-AUC ROC-AUC PR-AUC ROC-AUC PR-AUC ROC-AUC PR-AUC
Baseline TrialBench 0.475±0.027 0.255±0.006 0.569±0.010 0.295±0.008 0.550±0.012 0.279±0.008 0.477±0.022 0.256±0.007
HINT 0.540±0.022 0.272±0.009 0.535±0.009 0.267±0.005 0.474±0.019 0.251±0.006 0.548±0.021 0.273±0.014
ClinicalReTrial Logistic Reg. 0.606±0.019 0.326±0.015 0.583±0.011 0.306±0.008 0.621±0.016 0.350±0.017 0.550±0.022 0.280±0.010
GBDT 0.633±0.016 0.344±0.016 0.662±0.011 0.382±0.011 0.669±0.017 0.412±0.019 0.543±0.025 0.282±0.012
Table 4: Performance of multi-class clinical trial outcome prediction across trial phases.
Feature Importance Analysis.
We studied feature importance analysis with SHAP (lundberg2017shap). Consistent with our hypothesis, sentence-level eligibility, Drug-disease interaction features and endpoint alignment features are most important for outcomes prediction (Appendix A.1.3).
4.3 Agent’s Protocol Optimization
Having confirmed the simulation environment’s reliability, we evaluate ClinicalReTrialAgent’s ability to redesign failed clinical trials.
4.3.1 Convergence Analysis
Table 5 reports comprehensive convergence statistics across all trials. The Agent had 83.3% of protocol designs improved (50/60 successfully processed trials showed positive $\Delta p$), with 4 trials (6.7%) encountering agent failures where the system identified zero opportunities of potential redesign, occurring in the efficacy failure mode.
Table 5: Convergence analysis across 60 test trials.
Mode Trials Failures Pos. $\Delta p$ Threshold
Enrollment 20 0 20/20 (100%) 8/20 (40%)
Safety 20 0 18/20 (90%) 4/20 (20%)
Efficacy 20 4 12/20 (60%) 10/20 (50%)
Overall 60 4 (6.7%) 50/60 (83.3%) 22/60 (36.7%)
The system demonstrated efficient convergence patterns, with 15% (9/60) of trials exhibiting natural termination before iteration 5 due to exhausted modification space. Most trials used all 5 iterations, suggesting adaptive stopping could improve efficiency.
4.3.2 Trial Redesign Learning Dynamics
We examine the learning trajectory of successfully processed trials (56/60). The system achieved mean improvement of $\Delta p=+0.057$ (95% CI: [+0.040, +0.074]). Table 6 stratifies results by failure mode, while Figure 2 illustrates learning dynamics by iteration.
Table 6: Probability shift ($\Delta p$) analysis by failure mode across 56 trials. IQR is 25th-75th percentile range.
Mode Trials $p_{0}$ / $p_{final}$ $\Delta p$ IQR
Enrollment 20 0.506 / 0.563 +0.058 [+0.034, +0.068]
Safety 20 0.791 / 0.831 +0.070 [+0.032, +0.092]
Efficacy 16 0.813 / 0.859 +0.040 [+0.013, +0.039]
Overall 56 0.695 / 0.730 +0.057 [+0.029, +0.073]
Figure 2: Iterative learning dynamics stratified by failure mode. IQR bars span the 25th–75th percentiles (interquartile range) across runs.
Performance heterogeneity across failure modes reflects the differential amenability of clinical trial design elements to protocol-level intervention. Safety failures exhibit the largest improvements (mean $\Delta p=+0.070$, IQR [+0.032, +0.092]), as adverse events often stem from identifiable contraindication patterns that can be systematically addressed through eligibility refinement and dosage adjustment. Enrollment failures show substantial gains (mean $\Delta p=+0.058$), consistent with the observation that recruitment barriers frequently arise from overly restrictive or poorly specified inclusion criteria rather than fundamental feasibility constraints. Efficacy failures demonstrate the smallest yet statistically meaningful improvements (mean $\Delta p=+0.040$), as therapeutic effectiveness depends heavily on drug-disease compatibility, where sometimes protocol modifications alone cannot fix the essential drug failure.
As shown in Figure 2, the learning trajectory reveals major initial gains followed by decreasing returns. This pattern validates the knowledge distillation mechanism: high-quality modifications are identified early through rapid retrieval augmented analysis, while later iterations exploit narrower optimization opportunities by refining secondary parameters or addressing edge case contraindications.
4.3.3 Computational Efficiency
[…]
Table 7: Computational efficiency by failure mode.
[…]
Enrollment 20 0.171 +0.055 4.76
Safety 20 0.100 +0.054 3.00
Efficacy 16 0.088 +0.038 3.05
Overall 56 0.122 +0.053 3.68
4.4 Ablation Study on Self-improvement
To validate architectural components, we conducted paired ablation across 10 enrollment failure trials, systematically removing: (1) memory-guided iterative learning and (2) redesign pool optimization. Each trial was evaluated under all three conditions with identical initialization, enabling within-subjects comparison.
Both components contribute significantly and independently (Figure 3, Table 8). Memory removal degraded performance at iteration 1 ($\Delta p=+0.0131$, $p=0.042$) and iteration 5 ($\Delta p=+0.0190$), demonstrating immediate warm start benefits that strengthen over iterations. Removing the redesign pool yielded comparable degradation ($\Delta p=+0.0126$). These findings validate our design: memory enables the agent to manage intelligent exploration, while the redesign pool enables exploitation by reusing successful modifications.
Figure 3: Self-Improving ablation study across 10 trials. Full system (blue) outperforms setups without memory (red) or redesign pool optimization (green). Memory provides early benefits, and pool effects compound over iterations. IQR bars span the 25th–75th percentiles across runs.
Figure 4: ClinicalReTrial Agent’s flowchart on Poor Enrollment failed trial case study (NCT01298752, 2011-02-16), together with the real-world redesign (NCT01591161, 2012-05-02). The Agent’s iterative refinement of failure analysis and according modifications are demonstrated.
Table 8: Self-Improving ablation study. Stats computed using paired $t$-tests with $n=10$ trials. *$p$-value <0.05<0.05, **$p$-value <0.01<0.01.
Removed Iteration Ablation Full System $\Delta p$ $p$-value Cohen’s $d_{z}$
Memory 1 $024\pm 0.016 0.0$$37\pm 0.020 +0.0$13+0.013 $0.042^{*}$ $0.81$
Memory 5 $038\pm 0.016 0.0$$57\pm 0.023 +0.0$19+0.019 $0.007^{**}$ $1.10$
Pool 5 $045\pm 0.022 0.0$$57\pm 0.023 +0.0$13+0.013 $0.005^{**}$ $1.18$
4.5 Retrospective Case Studies: Real-World Validation
To validate ClinicalReTrial’s clinical applicability, we analyze trial pairs, where investigators successfully redesigned and re-executed failed protocols spanning enrollment, safety, and efficacy failure modes (Appendix A.5). Validating agent reasoning against real-world protocol redesigns provides critical insight into clinical applicability.
Here we present a poor enrollment redesign case (safety and efficacy cases in Appendix A.5): NCT01298752, a Phase 3 trial of Mapracorat (anti-inflammatory ophthalmic suspension) for post-cataract surgery inflammation that failed due to slow enrollment. Sponsored by Bausch & Lomb, the trial was subsequently redesigned and successfully executed as NCT01591161. Figure 4 illustrates ClinicalReTrial’s iterative refinement process across three optimization cycles. The Enrollment barrier (cataract surgery waiting requirement) is efficiently identified with positive reward provided by simulation environment. The agent also progressively explores the modification space: baseline AC cell requirements is successfully added as an enrichment criterion; while agent also explores the safety enhancement, but end up failing to align with real-world redesign.
5 Conclusion
In this work, we have presented ClinicalReTrial, a novel self-evolving agent framework that moves beyond passive clinical trial outcome prediction to enable proactive optimization of clinical trial protocols. By integrating an interpretable, accurate simulator with an autonomous Agent capable of causal reasoning and iterative design refinement, our system not only able to forecast trial success but also generates actionable modifications, including adjusting eligibility criteria, dosing regimens, or endpoint definitions to enhance feasibility and likelihood of success. Evaluated on standardized TrialBench benchmark, ClinicalReTrial achieves strong predictive performance while demonstrating the ability to discover significant protocol improvements with clinical best practices.
6 Limitations
Our framework has several limitations that suggest directions for future research. First, the simulation environment’s predictive accuracy creates potential for false improvement signals and missed opportunities, though validation filtering and iterative refinement provide partial mitigation; this constraint could be addressed by integrating future state-of-the-art prediction models as drop-in replacements for the GBDT oracle. Second, the system lacks adaptive convergence detection; 83.9% of trials exhausted the full 5-iteration budget rather than stopping when modification potential plateaus, suggesting the need for learned stopping criteria based on diminishing returns patterns. Third, retrospective case study analysis reveals tactical domain knowledge gaps: while the system excels at strategic-level reasoning, it may struggle with operational specifics such as selecting appropriate biomarkers, anticipating implementation constraints, and distinguishing when radical simplification outperforms incremental fortification, often over-relying on complexity multiplication where parsimony proves more effective. Future work should prioritize prospective validation in collaboration with clinical trial sponsors, integration of specialized biomarker knowledge bases to address tactical gaps, development of adaptive stopping mechanisms to improve computational efficiency, and expansion to larger-scale evaluation encompassing broader disease areas and trial designs.
Appendix A Appendix
A.1 Simulation Environment Details
A.1.1 Dataset Statistics
Dataset we used to train the prediction models comprises 20,769 clinical trials from TrialBench’s failure reason dataset. Table 9 shows the label distribution across four categories.
Table 9: Distribution of failure reason labels in the dataset.
Failure Reason Count Percentage
Success 9,939 47.8%
Poor Enrollment 7,229 34.8%
Inefficacy 2,217 10.7%
Adverse Effect 1,384 6.7%
Total 20,769 100.0%
The class imbalance reflects real-world trial outcomes: enrollment challenges are the most common failure mode (34.8%), followed by efficacy gaps (10.7%), while safety failures are relatively rare (6.7%) due to rigorous preclinical screening. Success cases (47.8%) include trials that completed without major protocol violations or early termination.
Due to computational cost constraints, we randomly select a stratified sample of 60 trials from the test set (20 enrollment, 20 safety, 20 efficacy), ensuring representation across failure modes and trial phases. Table 10 presents the phase composition.
Table 10: Test distribution by trial phase.
Phase Count %
Phase 1 8 13.3
Phase 2 27 45.0
Phase 3 15 25.0
Phase 4 10 16.7
Total 60 100
A.1.2 Encode Details
This appendix provides comprehensive implementation details for the Simulation Environment described in §3.2.4, including encoder pretraining procedures, model training hyperparameters, and detailed validation results.
Text Features.
Textual contents are encoded using BioBERT (lee2020biobert), a domain-adapted language model pre-trained on PubMed abstracts and PMC full-text articles. Critically, we diverge from prior work by decomposing eligibility criteria at the sentence level rather than treating them as monolithic text blocks. For each text field $\mathcal{T}$, we decompose it into sentences $\mathcal{T}=\{s_{1},s_{2},\ldots,s_{n}\}$. Each sentence is encoded via BioBERT and the final text embedding is obtained via max pooling:
$\mathbf{e}_{s_{i}}=\text{BioBERT}(s_{i}),\quad\mathbf{h}_{\mathcal{T}}=\max_{i=1}^{n}\mathbf{e}_{s_{i}}$ (2)
This sentence-level representation preserves granularity essential for aspect-specific modification—ClinicalReTrialAgent can target individual criteria rather than generic protocol summaries.
Graph Features.
We incorporate pre-trained molecular and disease encodings to capture pharmacological properties and disease characteristics.
Drug Molecular Graphs. Each drug molecule $m$ is represented as a graph $\mathcal{G}_{m}=(\mathcal{V},\mathcal{E})$ where nodes $v\in\mathcal{V}$ are atoms and edges $(u,v)\in\mathcal{E}$ are bonds. We employ Message Passing Neural Networks (MPNNs) to aggregate neighborhood information over $L$ iterations:
$\mathbf{m}_{uv}^{(l)}=\text{ReLU}\left(W_{i}\cdot[\mathbf{f}_{u}\oplus\mathbf{f}_{uv}]+W_{h}\cdot\sum_{w\in\mathcal{N}(u)\setminus v}\mathbf{m}_{wu}^{(l-1)}\right),$ (3)
where $\mathbf{m}_{uv}^{(l)}\in\mathbb{R}^{d_{\text{mpnn}}}$ is the message from atom $u$ to atom $v$ at layer $l$, $\mathcal{N}(u)$ denotes neighbors of $u$, $\oplus$ denotes concatenation, and $W_{i},W_{h}$ are learnable transformation matrices. After $L$ message passing iterations, node embeddings are computed as:
$\mathbf{h}_{u}=\text{ReLU}\left(W_{o}\cdot\left[\mathbf{f}_{u}\oplus\sum_{v\in\mathcal{N}(u)}\mathbf{m}_{vu}^{(L)}\right]\right).$ (4)
The graph-level drug embedding is obtained via global average pooling:
$\mathbf{h}_{\text{drug}}=\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}\mathbf{h}_{u}\in\mathbb{R}^{d_{\text{mpnn}}}$ (5)
For trials with multiple drugs, we average their embeddings. The MPNN encoder is pretrained on pharmacokinetic (ADMET) tasks, then fine-tuned on trial outcome labels (details in Appendix A.1).
Disease Hierarchical Encoding. Each disease is represented by an ICD-10 code $d_{i}$ following a hierarchical taxonomy with ancestors $\mathcal{A}(d_{i})=\{a_{1},a_{2},\ldots,a_{p}\}$. We use Graph-based Attention Model (GRAM) to encode hierarchical disease information. Each code $c$ has a learnable base embedding $\mathbf{e}_{c}\in\mathbb{R}^{d_{\text{gram}}}$. The hierarchical embedding for disease $d_{i}$ is computed as an attention-weighted sum over itself and its ancestors:
$\mathbf{h}_{d_{i}}=\sum_{a_{j}\in\mathcal{A}(d_{i})\cup\{d_{i}\}}\alpha_{ji}\cdot\mathbf{e}_{a_{j}}$ (6)
where the attention weight $\alpha_{ji}$ measures the relevance of ancestor $a_{j}$ to the current disease $d_{i}$:
$\alpha_{ji}=\frac{\exp(\phi([\mathbf{e}_{a_{j}}\oplus\mathbf{e}_{d_{i}}]))}{\sum_{a_{k}\in\mathcal{A}(d_{i})\cup\{d_{i}\}}\exp(\phi([\mathbf{e}_{a_{k}}\oplus\mathbf{e}_{d_{i}}]))}$ (7)
where $\phi(\cdot):\mathbb{R}^{2d_{\text{gram}}}\rightarrow\mathbb{R}$ is a learnable single-layer network. For trials targeting multiple diseases, we average their embeddings. The GRAM encoder is initialized with the ICD-10 hierarchical ontology, then fine-tuned on historical trial success rates (details in Appendix A.1).
Tabular Features.
We encode structured trial metadata through a modular pipeline that processes categorical attributes, demographic constraints, administrative properties, and enrollment characteristics. The pipeline extracts 29 numerical features.
Problem Formulation and Dataset. We formulate clinical trial outcome prediction as a binary classification problem over three distinct failure modes: poor enrollment, safety/drug adverse effect, and drug inefficacy. We train models separately for each failure mode, enabling ClinicalReTrialAgent to target specific causes during protocol optimization. Our experiments utilize the TrialBench dataset (chen2025trialbench), which contains over 12,000 annotated clinical trials spanning Phase I through Phase IV, with each trial labeled according to outcome. The dataset provides multi-modal features including drug molecular structures, disease ICD-10 codes, eligibility criteria text, trial metadata, and intervention details. Following standard practice to avoid temporal leakage, we partition data chronologically by trial completion year. According to Table 11, features are encoded into total dim=6,173.
Table 11: Feature specification summary. Novel contributions include sentence-level eligibility parsing and fine-tuned molecular-disease encoders.
Category Component Dim Method Novel
Text Study Design 768 BioBERT
Dosage 768 BioBERT
Intervention 768 BioBERT + pooling
Condition 768 BioBERT + pooling
Eligibility Inclusion 768 BioBERT + pooling
Eligibility Exclusion 768 BioBERT + pooling
Graph Drug (ADMET) 768 MPNN (fine-tuned)
Disease (ICD) 768 GRAM (fine-tuned)
Tabular Categorical Features 18 One-Hot
Age constraints 2 Unit normalization
Multi-hot indicators 9 Binary encoding
Total 6,173
Feature Concatenation and Prediction. All feature modalities are concatenated into a single input vector:
$\mathbf{x}_{\text{trial}}=[\mathbf{h}_{\text{design}};\mathbf{h}_{\text{dose}};\mathbf{h}_{\text{interv}};\mathbf{h}_{\text{cond}};\mathbf{h}_{\text{incl}};\mathbf{h}_{\text{excl}};\mathbf{h}_{\text{drug}};\mathbf{h}_{\text{disease}};\mathbf{f}_{\text{tabular}}]\in\mathbb{R}^{6173}$ (8)
where semicolons denote concatenation. For each failure mode $\tau\in\{\text{enrollment, safety, efficacy}\}$, we train a separate LightGBM classifier $\mathcal{M}_{\tau}$ that predicts trial success probability, optimizing binary cross-entropy loss. The predicted probability $\hat{y}=\mathcal{M}_{\tau}(\mathbf{x}_{\text{trial}})\in[0,1]$ serves as the reward signal for evaluating protocol modifications in the agent system.
Model Training and Validation. We employ LightGBM (ke2017lightgbm) for its computational efficiency with high-dimensional sparse features. Three independent models are trained for enrollment, safety, and efficacy failure prediction using cross-validation with early stopping. The trained GBDT models achieve strong predictive performance across all failure modes (PR-AUC >> 0.75) with well-calibrated probability estimates, validating the simulation environment as a reliable proxy for real trial outcomes.
A.1.3 Ablation Study
Word-Level Attention Analysis. Figure 5 demonstrates the word-level attention weights captured by BioBERT embeddings in the TrialDura model, visualized through Shapley values. The heatmap reveals that clinical keywords such as “woman,” “contraception,” receive the highest attention weights (0.0208–0.0274), while functional words like prepositions and conjunctions are assigned lower weights. This attention distribution indicates that the model effectively focuses on medically relevant terminology when processing eligibility criteria, suggesting that domain specific language models can automatically identify critical phrases without explicit feature engineering.
- potentially fertile woman without $\beta$-hcg
0.0001 0.0075 0.0150 0.0208 0.0001 0.0165
negative harvested until 48 hours before
0.0179 0.0175 0.0144 0.0142 0.0157 0.0153
operation or not using acceptable contraception
0.0184 0.0177 0.0168 0.0163 0.0156 0.0212
for participation in this study
0.0245 0.0166 0.0088 0.0155 0.0274
Figure 5: Visualization of text segments in the BioBERT encoder’s output, illustrating Shapley values derived from Clinical Trials. Shapley values correspond to attention weights, with darker colors indicating higher weights.
Sentence-Level Eligibility Weights. Table 6 illustrates a example of sentence-level importance scores within the inclusion criteria for trial NCT01102504, normalized across all eligibility statements, with weighted importance calculated on predict probability shift if masking out each eligibility protocols. The model assigns highest weights (0.20–0.25) to sentences describing acute cerebrovascular events such as “Transient ischemic attack (TIA)” and “Stroke (ipsilaterally to the stenotic artery),” while demographic criteria like age receive minimal attention (0.07). Notably, the quantitative stenosis threshold $30\%$ stenosis on initial B-mode ultrasonography imaging” receives substantial weight (0.18), indicating that the model prioritizes disease severity markers and clinical events over basic demographic qualifications when predicting trial outcomes.
Table 12: Inclusion Criteria with Sentence Importance (Color-coded)
NCT01102504 Eligibility Criteria Protocols Weight
Inclusion Criteria: 0.03
 - Age 40–90 years old, 0.07
 - Clinically documented carotid symptomatic atherosclerotic disease (symptomatic disease will be considered if one of the following has occurred within 2 months prior to symptoms:) 0.12
   1. Amaurosis fugax 0.10
   2. Transient ischemic attack (TIA) 0.20
   3. Stroke (ipsilaterally to the stenotic artery) 0.25
- $30\%$ stenosis on initial B-mode ultrasonography imaging, 0.18
 - Written, informed consent. 0.05
Encodes Contributions Revealed Through Ablation Analysis. Figure 6 presents the relative importance of different encoders across three prediction tasks through systematic masking experiments. By individually masking each encoder and measuring the resulting PR-AUC drop, we quantify each component’s contribution to enrollment, safety, and efficacy outcome predictions. The analysis reveals task-specific dependency patterns: certain encoders prove critical for particular outcomes, with their removal causing substantial performance degradation, while showing minimal impact on other tasks. This heterogeneous importance distribution demonstrates that different aspects of trial design and patient characteristics drive distinct clinical endpoints. The varying magnitudes of PR-AUC drops across tasks validate the multi-task learning framework’s ability to capture task-specific representations while identifying which shared features are most crucial for each prediction objective.
Figure 6: Feature importance of each encoder on 3 classification tasks (enrollment, safety, and efficacy), measured by PR-AUC drop when the encoder is masked out during prediction.
A.2 Analysis Agent Details
The Analysis Agent implements a domain-aware ReAct reasoning pipeline adapted by failure mode (enrollment, safety, efficacy). Novel components include adverse event profiling, statistical power assessment, and design-level pivots. Table 13 summarizes failure-mode-specific adaptations.
Table 13: Analysis Agent variants by failure mode.
Component Enrollment Safety Efficacy
Profiling None Adverse Event Profiling (severity, organ systems, root cause) Efficacy Gap Profiling (observed vs expected)
Classification 4 categories 5 categories (adds safety_inadequate) 4 categories (weights enrichment higher)
Assessments None Dosage + AE profile Dosage + Outcome + Power analysis
Prioritization Confidence-based Safety-first (toxicity reduction priority) Simplicity-first tiered (PRIMARY/SECONDARY/TERTIARY)
Protocol Classification Role: Clinical researcher classifying eligibility criteria Task: Classify criteria into 4 categories with confidence scores [0-1] Context:
Phase: Phase 2
Mechanism: X inhibits Y pathway
Endpoint: Measuring Z at 12 weeks
Criteria to Classify:
<criterion aspect_name="eligibility/inclusion_criteria" index="1">
Must wait for fellow eye surgery until study completion
</criterion>
<criterion aspect_name="eligibility/exclusion_criteria" index="2">
Any prior participation in drug trials within 12 months
</criterion>
Categories: 1. PARTICIPATION_BARRIER: Timing/waiting requirements, administrative hurdles 2. SAFETY_EXCLUSION: Medical risks (allergies, drug interactions, severe conditions) 3. SELECTION_CRITERION: Defines WHO is eligible (disease type, procedure type, demographics) 4. ENRICHMENT_CRITERION: Selects likely responders (biomarkers, mechanism-aligned traits) For each criterion, assign scores [0-1] to ALL categories, pick PRIMARY (highest), give 1-sentence reason. Output Format:
<classification aspect_name="eligibility/inclusion_criteria" index="1">
<participation_barrier_score>0.92</participation_barrier_score>
<safety_exclusion_score>0.05</safety_exclusion_score>
<selection_criterion_score>0.20</selection_criterion_score>
<enrichment_criterion_score>0.10</enrichment_criterion_score>
<primary_category>PARTICIPATION_BARRIER</primary_category>
<reasoning>Waiting requirement for fellow eye surgery is a strong
participation barrier with no medical justification.</reasoning>
</classification>
Mechanism Alignment Check Role: Clinical researcher evaluating mechanism alignment Task: Check if criteria select mechanism-appropriate patients and detect missing enrichment Questions: 1. Do we select patients who HAVE the target condition this mechanism treats? 2. Do we select patients with baseline values allowing measurement of endpoint Y? 3. Are safety exclusions too broad, blocking potential responders? If missing enrichment (no criteria selecting treatment-responsive patients): Propose ONE objective criterion with: measurement method, threshold, timing Must be measurable (grades/scores/labs), not subjective ("anticipated"/"likely") Output Format:
<mechanism_analysis>
Current criteria define cataract surgery candidates but lack enrichment
for inflammation severity. Waiting requirement blocks eligible patients
without medical benefit.
</mechanism_analysis>
<missing_enrichment_criterion>
Add inclusion: Baseline anterior chamber cell grade $\geq$$2+ (SUN criteria)
measured within 7 days of enrollment. Selects patients with measurable
inflammation for mechanism-aligned response assessment.
</missing_enrichment_criterion>
Adverse Event Profiling Role: Clinical researcher analyzing safety failures Task: Parse and categorize adverse events for safety redesign Input:
Adverse events: Hepatotoxicity (Grade 3, 25%), elevated AST/ALT (Grade 2, 40%)
Intervention: Drug X (oral, 100mg daily for 28 days)
Mechanism: Inhibits enzyme Y in Z pathway
Instructions: 1. SEVERITY CLASSIFICATION: Extract Grade 3-5 events (dose-limiting), Grade 2 (tolerability) 2. ORGAN SYSTEM MAPPING: Map toxicity to organ (Liver, Kidney, Bone marrow, Heart, GI) 3. MECHANISM CONSISTENCY: Does toxicity match expected mechanism? 4. DOSE-RESPONSE INFERENCE: Dose-dependent? Acute or cumulative? 5. PRIORITY RANKING: CRITICAL (Grade 3+ >10%), HIGH (Grade 2+ >30% OR any Grade 4+) 6. ROOT CAUSE HYPOTHESIS: Excessive dose, inadequate exclusions, off-target effects? Output Format:
<adverse_event_profile>
<primary_toxicity>
  <event>Hepatotoxicity</event>
  <grade>3</grade>
  <incidence>25%</incidence>
  <organ_system>Liver</organ_system>
  <priority>CRITICAL</priority>
  <dose_dependent>likely</dose_dependent>
</primary_toxicity>
<mechanism_consistency>
UNEXPECTED - mechanism does not predict liver toxicity
</mechanism_consistency>
<root_cause_hypothesis>
Likely excessive dose (100mg exceeds typical range) or missing hepatic
impairment exclusion. Drug metabolism may saturate at high doses.
</root_cause_hypothesis>
<critical_gaps>
<gap>Exclude patients with baseline AST/ALT >2x ULN</gap>
<gap>Exclude patients with Child-Pugh Class B or C cirrhosis</gap>
</critical_gaps>
</adverse_event_profile>
Design-Level Pivots Role: Clinical trial designer proposing trial-level redesign Task: Propose high-level trial redesign (not just criteria tweaks) Context:
Phase: Phase 2
Mechanism: Inhibits enzyme Y
Failure: Grade 3 hepatotoxicity 25%
Redesign archetype: PK_SAFETY_FOLLOWUP
Primary outcome: Safety assessment at 28 days
Dosage assessment: EXCESSIVE (100mg daily exceeds safe exposure)
Design Pivot Rules: If archetype is PK_SAFETY_FOLLOWUP or main failure is safety-driven: Prefer PK_SAFETY or DOSE_FINDING trial type Prefer PK-focused primary endpoints Prefer simpler care model with lower background risk Prefer simpler dosing (single-dose or short-duration) When systemic toxicity suspected: Consider more local/regional route to reduce systemic exposure Consider smaller, denser design (PK_SINGLE_ARM with intensive sampling) Output Format:
<design_pivots>
<trial_type>PK_SAFETY</trial_type>
<endpoint_family>PK_SAFETY</endpoint_family>
<dose_regimen_direction>SIMPLER</dose_regimen_direction>
<route_change>CONSIDER_ALTERNATIVE_ROUTE</route_change>
<proposed_route>Consider single 25mg dose with intensive PK sampling
over 7 days, or switch to subcutaneous administration to reduce
first-pass hepatic metabolism</proposed_route>
<sample_size_direction>SMALLER</sample_size_direction>
<design_structure>PK_DOSE_FINDING</design_structure>
<proposed_primary_outcome>Area under curve (AUC) and peak liver enzyme
elevation (AST/ALT) at 24h, 48h, 72h post-dose</proposed_primary_outcome>
<summary>Pivot from Phase 2 efficacy trial to Phase 1b/2a PK safety
study. Reduce dose to 25mg single administration with intensive PK and
liver function monitoring. Alternative route (subcutaneous) may bypass
hepatic first-pass effect. Smaller sample (N=20-30) adequate for PK
characterization. Expected to reduce Grade 3+ hepatotoxicity from 25%
to <5%.</summary>
</design_pivots>
Trade-off Analysis Role: Clinical pharmacologist analyzing dosage for EFFICACY failure Task: Analyze DOSAGE trade-offs (not safety) Context:
Current dosage: 50mg oral daily for 21 days
Dosage assessment: SUBOPTIMAL
  Classification reasoning: Phase 1 MTD was 100mg daily. Current 50mg
  dose is at 50% of MTD with acceptable safety. PK data shows linear
  dose-response up to 80mg.
  Suggested: Escalate to 75mg daily
Mechanism: Inhibits receptor X
Efficacy Gap: ORR 15% vs 30% (gap: 15%)
Power Assessment: LIKELY underpowered, Root Cause: BOTH
Instructions: 1. RECOMMENDATION: MODIFY (escalate) or KEEP (defer) 2. IMPACTS: efficacy_signal [++], enrollment [0], safety [-], mechanism [ALIGNED] 3. CONFIDENCE: High (0.80-0.90) if clear PK/PD data 4. REASONING: Include feasibility (Time: X-Ymo; Burden: LOW|MED|HIGH; Cost: Zx) Output Format:
<dosage_tradeoff>
<recommendation>MODIFY</recommendation>
<efficacy_signal>++</efficacy_signal>
<enrollment>0</enrollment>
<safety>-</safety>
<mechanism_alignment>ALIGNED</mechanism_alignment>
<confidence>0.85</confidence>
<reasoning>Escalating to 75mg (75% of MTD) expected to improve ORR by
10-15 percentage points based on linear PK and Phase 1 exposure-response.
Safety risk manageable (Grade 2 toxicity may increase from 20% to 30%).
FEASIBILITY: Time: 1-3mo; Burden: LOW; Cost: 1.2x (simple dose
adjustment, no formulation change).</reasoning>
</dosage_tradeoff>
A.3 Augmentation Agent Details
Table 14: Augmentation Agent novel features by failure mode. Process differ a little in dosage modification strategy across failure modes.
Feature Enrollment Safety Efficacy
Dosage Strategy N/A Reduce ($\downarrow$25-50%, fractionation, pulse) Escalate ($\uparrow$25-50%, loading, dose-dense)
Outcome Strategy N/A Add safety qualifications Switch to feasible endpoint
Domain Focus Enrich participation Tighten safety exclusions Add biomarker enrichment
A.3.1 Few-Shot Learning Mechanism
Few-Shot Example Injection (Shared Structure) Matching Logic:
# LIST aspects (eligibility criteria)
prev_rules["seen_indices"][aspect_name][str(aspect_index)]
# STRING aspects (dosage, target_primary_outcome)
prev_rules["seen_indices"][aspect_name]["None"]
Injected Section in MODIFY Prompts (Iteration 2+):
<few_shot_examples>
Previous iteration examples for THIS EXACT criterion:
EXCELLENT:
  - [Example that led to excellent validation score]
  - [Another excellent example]
GOOD:
  - [Example that led to good validation score]
MODERATE:
  - [Example with moderate validation score]
BAD:
  - [Example that validation agent rejected]
BANNED:
  - [Example that was explicitly banned (safety violation)]
Generate variations that learn from EXCELLENT/GOOD patterns,
avoid BAD patterns, and NEVER replicate BANNED augmentations.
</few_shot_examples>
Effect: LLM learns from previous iteration’s successes/failures. Only available iteration 2+ after prev_rules established.
A.3.2 Augmentation Prompts by Failure Mode
Eligibility Example Role: Clinical researcher generating criterion variations Task: Generate num_augment variations with few-shot guidance Input:
Original criterion: "Must wait for fellow eye surgery until completion"
Strategy: "Delete waiting requirement to increase enrollment"
Failure mode: Enrollment
Adaptive num_augment: 3 (medium variance)
Few-Shot Examples (if iteration 2+):
EXCELLENT: "No waiting period required between surgeries"
GOOD: "Fellow eye surgery allowed concurrent with study"
BAD: "Reduced wait from 6 months to 3 months" (still a barrier)
BANNED: "Must complete fellow eye surgery before enrollment" (contradicts)
Universal Requirements: Each variation MUST directly implement the Strategy Preserve clinical intent, make more operational/measurable/specific Objective and quantifiable (use thresholds, time windows, methods) Avoid vague language: "anticipated", "expected", "likely", "may", "severe" Maintain consistency with safety and mechanism of action All variations distinct from each other Output:
<augmentations>
<augmentation>No waiting period required between fellow eye surgeries</augmentation>
<augmentation>Fellow eye surgery allowed at any time during study</augmentation>
<augmentation>Bilateral surgery candidates eligible without delay</augmentation>
</augmentations>
Dosage Example Role: Clinical pharmacologist reducing dosage to minimize toxicity Task: Generate num_augment dosage reductions Input:
Original dosage: 100mg oral daily for 28 days
Adverse events: Hepatotoxicity (Grade 3, 25%), AST/ALT elevation (Grade 2, 40%)
Strategy: Reduce dose to decrease Grade 3+ hepatotoxicity to <10%
Adaptive num_augment: 5 (high variance)
Few-Shot Examples (iteration 3):
EXCELLENT: "50mg oral daily (50% reduction, expected toxicity <8%)"
GOOD: "50mg BID (fractionated, reduces Cmax and hepatic load)"
MODERATE: "75mg oral daily (25% reduction, may be insufficient)"
BAD: "90mg oral daily (only 10% reduction)"
BANNED: "100mg every other day (same cumulative exposure)"
[…]
<augmentations>
<augmentation>
<dosage_modification>50mg oral daily for 28 days</dosage_modification>
<rationale>50% dose reduction expected to reduce hepatotoxicity
from 25% to <8% based on linear dose-toxicity relationship</rationale>
</augmentation>
<augmentation>
<dosage_modification>40mg BID (total 80mg daily, fractionated)</dosage_modification>
<rationale>Fractionated dosing reduces Cmax by ˜40%, lowering peak
hepatic exposure while maintaining 80% dose intensity</rationale>
</augmentation>
<augmentation>
<dosage_modification>50mg on days 1-5, off days 6-7 each week</dosage_modification>
<rationale>Pulse dosing (71% intensity) allows hepatic recovery,
expected to reduce Grade 3+ events to <10%</rationale>
</augmentation>
</augmentations>
A.4 Agent Output Template
This section presents the structured output format produced by the agent pipeline. The complete output is stored as JSON and includes trial data, ReAct reasoning traces, and generated protocol modifications.
Agent Pipeline Output Structure (Generic Template)
  "trial_data": {
    "nct_id": "NCT########",
    "phase": "Phase X",
    "condition": "[Disease/Condition]",
    "intervention/intervention_name": "[Intervention Name]",
    "failure_reason": "[enrollment|safety|efficacy]",
    "adverse_events": "[Adverse event summary or ’Not specified’]",
    "eligibility/inclusion_criteria": [
      "[Inclusion criterion 1]",
      "[Inclusion criterion 2]",
    "eligibility/exclusion_criteria": [
      "[Exclusion criterion 1]",
      "[Exclusion criterion 2]",
    "dosage": "[Dosage regimen]",
    "target_primary_outcome": "[Primary outcome description]"
  "trial_context": {
    "phase": "Phase X",
    "mechanism_of_action": "[Mechanism description]",
    "primary_endpoint_type": "[Endpoint type description]",
    "redesign_archetype": "[PK_SAFETY_FOLLOWUP | DOSE_FINDING_REDESIGN |
                           ENRICHED_EFFICACY_RETRY | OTHER]",
    "index_surgical_model": "[Care/procedural model description]"
  "react_reasoning": {
    "step0_contextualize": {
      "phase": "Phase X",
      "mechanism_of_action": "[Mechanism extracted by LLM]",
      "adverse_event_profile": {
        "primary_toxicity": {
          "event": "[Primary adverse event]",
          "grade": "[0-5]",
          "incidence": "[X%]",
          "priority": "[CRITICAL|HIGH|MEDIUM|LOW]"
        "root_cause_hypothesis": "[Root cause analysis by LLM]"
      "dosage_assessment": {
        "classification": "[EXCESSIVE|BORDERLINE|APPROPRIATE|SUBOPTIMAL]",
        "reasoning": "[Dosage assessment reasoning]"
    "step1_classification": [
        "aspect_name": "eligibility/[inclusion|exclusion]_criteria",
        "aspect_index": N,
        "criterion_text": "[Original criterion text]",
        "participation_barrier_score": 0.X,
        "safety_exclusion_score": 0.X,
        "selection_criterion_score": 0.X,
        "enrichment_criterion_score": 0.X,
        "primary_category": "[PARTICIPATION_BARRIER | SAFETY_EXCLUSION |
                            SELECTION_CRITERION | ENRICHMENT_CRITERION]",
        "reasoning": "[Classification reasoning]"
        "aspect_name": "eligibility/[inclusion|exclusion]_criteria",
        "aspect_index": M,
        "criterion_text": "[Original criterion text]",
        "primary_category": "[Category]",
        "reasoning": "[Classification reasoning]"
    "step2_mechanism_alignment": "[3-4 sentences on whether existing
                                  criteria + dosage maximize success
                                  probability for this failure mode]",
    "step3_tradeoff_analysis": [
        "aspect_name": "eligibility/[inclusion|exclusion]_criteria",
        "aspect_index": N,
        "enrollment_impact": "[--|-|0|+|++]",
        "efficacy_signal_impact": "[--|-|0|+|++]",
        "safety_risk_impact": "[--|-|0|+|++]",
        "mechanism_alignment": "[ESSENTIAL|ALIGNED|NEUTRAL|MISALIGNED]",
        "net_recommendation": "[KEEP|MODIFY|DELETE|ADD]",
        "confidence": 0.XX,
        "reasoning": "[Trade-off reasoning with feasibility encoding]"
        "aspect_name": "[dosage|target_primary_outcome|surgical_model|...]",
        "aspect_index": null,
        "enrollment_impact": "[Impact symbol]",
        "safety_risk_impact": "[Impact symbol]",
        "net_recommendation": "[MODIFY|ADD]",
        "confidence": 0.XX,
        "reasoning": "[Trade-off reasoning]"
    "step4_prioritization": "[6-8 sentences with tiered recommendations
                             (PRIMARY/SECONDARY/TERTIARY), timeline, and
                             confidence level]",
    "step5_synthesis": "[4-6 sentences synthesizing failure analysis with
                        quantification, expected benefits, trade-offs, and
                        overall confidence]"
  "aspect_li": [
      "aspect_name": "eligibility/[inclusion|exclusion]_criteria",
      "aspect_index": N,
      "original_value": "[Original criterion text]",
      "aspect_type": "list",
      "analysis": {
        "timestamp": "YYYY-$M$-DDTHH:$M$:SS",
        "failure_analysis": "[Analysis from step3 trade-off reasoning]",
        "impact_level": "[MAJOR|MINOR|NOT_RELATED]",
        "action_type": "[MODIFY|DELETE]",
        "strategy": "[Strategy from Analysis Agent]",
        "confidence": 0.XX
      "augment": {
        "timestamp": "YYYY-$M$-DDTHH:$M$:SS",
        "augment_val_li": [
          "[Augmentation 1]",
          "[Augmentation 2]",
          "[Augmentation 3]"
      "aspect_name": "eligibility/[inclusion|exclusion]_criteria",
      "aspect_index": null,
      "original_value": "N/A",
      "aspect_type": "list",
      "analysis": {
        "timestamp": "YYYY-$M$-DDTHH:$M$:SS",
        "failure_analysis": "[Analysis for ADD action]",
        "impact_level": "MAJOR",
        "action_type": "ADD",
        "strategy": "[Strategy from Analysis Agent]",
        "confidence": 0.XX
      "augment": {
        "timestamp": "YYYY-$M$-DDTHH:$M$:SS",
        "augment_val_li": [
          "[New criterion 1]",
          "[New criterion 2]",
          "[New criterion 3]"
      "aspect_name": "[dosage|target_primary_outcome]",
      "aspect_index": null,
      "original_value": "[Original value for string aspect]",
      "aspect_type": "string",
      "analysis": {
        "timestamp": "YYYY-$M$-DDTHH:$M$:SS",
        "failure_analysis": "[Analysis for string aspect]",
        "impact_level": "MAJOR",
        "action_type": "MODIFY",
        "strategy": "[Strategy from Analysis Agent]",
        "confidence": 0.XX
      "augment": {
        "timestamp": "YYYY-$M$-DDTHH:$M$:SS",
        "augment_val_li": [
          "[Modified value 1]",
          "[Modified value 2]",
          "[Modified value 3]"
A.5 Case Study Details
We validate ClinicalReTrialAgent’s reasoning against real-world protocol modifications provides critical insight into clinical applicability. We analyze three trial pairs where investigators redesigned and successfully re-executed failed protocols, enabling direct comparison between expert redesign decisions and ClinicalReTrialAgent’s proposals. Each case represents a distinct failure mode: NCT01298752 (poor enrollment), NCT01919190 (safety/adverse effects), and NCT02169336 (efficacy inadequacy).
Poor Enrollment.
To validate agent redesign quality against real-world outcomes, we analyze NCT01298752, a Phase 3 trial of Mapracorat (anti-inflammatory ophthalmic suspension) for post-cataract surgery inflammation that failed due to poor enrollment. Sponsored by Bausch & Lomb, the trial was subsequently redesigned and successfully executed as NCT01591161. Table 15 compares the real-world redesign with ClinicalReTrialAgent’s proposals.
Table 15: Agent-proposed modifications alignment check with real-world protocol redesign for poor enrollment, ClinicalReTrialAgent’s proposed modifications, categorizing alignment as: ✓(perfect match), $\sim$ (strategic alignment, tactical differences), or $\times$ (missed or incorrect).
Modification Type Real-World Redesign Agent Proposal Match Impact Level
Enrollment Barrier DELETE: "subjects must be willing to wait to undergo cataract surgery…" DELETE: "subjects must be willing to wait to undergo cataract surgery…" Major, removed primary barrier
Quality Enrichment ADDED: AC cells $\geq$ Grade 2 (6-15 cells) ADD: Require baseline AC cells $\geq$2 within 7 days Major, critical enrichment criteria
[…]
The primary enrollment barrier in the failed trial was a timing restriction requiring subjects to “wait to undergo cataract surgery on the fellow eye until after the study has been completed”—a constraint that excluded bilateral cataract patients unwilling or unable to delay their second surgery. Both the real-world redesign and ClinicalReTrialAgent correctly identified this as the critical obstacle and proposed its removal. Additionally, both approaches recognized the need for enrichment criteria: the real-world redesign added specific postoperative inflammation thresholds (AC cells $\geq$ Grade 2) to ensure enrolled patients exhibited measurable inflammation suitable for treatment evaluation, while ClinicalReTrialAgent proposed conceptually similar criteria targeting “mild to moderate inflammation”. However, the agent failed to capture domain-specific refinements present in the real-world redesign, including baseline safety standardization (requiring Grade 0 inflammation at screening) and operational clarity improvements (specifying exclusion of active external ocular disease). These tactical gaps highlight the agent’s limitations in translating strategic insights into clinically precise protocol language.
Safety/Adverse Events.
To validate agent redesign quality against real-world outcomes, we analyze NCT01919190, a Phase 4 trial of EXPAREL (liposomal bupivacaine) via TAP infiltration for post-surgical pain in lower abdominal procedures that failed due to severe adverse events (postoperative abdominal hemorrhage, 33.3% incidence). Sponsored by Pacira Pharmaceuticals, the drug was subsequently redesigned and successfully executed as NCT02199574 in a different surgical context. Table 16 compares the real-world redesign with ClinicalReTrialAgent’s proposals.
Table 16: Real-world validation (NCT01919190, Safety/Adverse Events): We compare the real-world changes with ClinicalReTrialAgent’s proposed modifications, categorizing alignment as: ✓(perfect match), $\sim$ (strategic alignment, tactical differences), or $\times$ (missed or incorrect).
Change Type Real-World Redesign ClinicalReTrialAgent Proposal Match Impact Level
Major Redesigns (Critical to Safety Success)
[…]
Dosage Reduction REDUCED by 50%: 266mg/20mL (60mL total volume) $\rightarrow$ 133mg/10mL (single dose, no dilution specified) REDUCED by $\sim$50%: Proposed 133mg in 20mL saline per validated option (total 40mL) Correct magnitude and direction
Surgical Model CHANGED procedure entirely: Lower abdominal surgeries (laparoscopic hysterectomy/myomectomy/colectomy with TAP infiltration) $\rightarrow$ Tonsillectomy (intraoperative infiltration to surgical site) Missing $\times$ Missing
Minor Refinements (Safety Improvements, Non-Critical to Success)
Eligibility Criteria SIMPLIFIED: Removed all TAP-specific anatomical exclusions, complex surgical requirements, chronic opioid exclusions, pain medication washout requirements, metastatic disease exclusions, substance abuse history exclusions; retained only: hypersensitivity to local anesthetics, investigational drug washout, pregnancy/nursing exclusions, and general “significant medical conditions” clause ADDED bleeding-specific exclusions: “Patients with history of bleeding disorders or on anticoagulant therapy” + liver dysfunction (Child-Pugh B/C) criteria; KEPT all 10 original complex exclusions including chronic opioid use, metastatic disease, substance abuse history, pain medication restrictions $\times$ Over-engineered restrictions vs. radical simplification
The primary safety issue in the failed trial was postoperative abdominal hemorrhage (33.3% incidence), attributed to excessive systemic exposure from high-volume TAP infiltration in hemorrhage-prone surgical sites. Both the real-world redesign and ClinicalReTrialAgent correctly identified the fundamental need to pivot from an efficacy trial to a PK/safety study and to reduce dosage by 50%, demonstrating strong diagnostic capability and appropriate dose-finding reasoning. However, the real-world approach implemented several structural changes largely absent from or contradicted by the agent’s proposal: radical surgical model change (lower abdominal surgeries $\rightarrow$ tonsillectomy), eliminating hemorrhage-prone anatomical sites entirely rather than attempting to “broaden” or “standardize” the same problematic surgical context; drastic scope reduction to a 12 patient PK characterization study rather than maintaining Phase 4 scale; and dramatic eligibility simplification, removing 6 of 10 complex exclusion criteria (chronic opioid use, metastatic disease, substance abuse, pain medication washout, TAP-specific anatomical concerns) to focus enrollment on the core safety profile.
Efficacy Inadequacy.
To validate agent redesign quality against real-world outcomes, we analyze NCT02169336, a Phase 2 trial of intranasal Dexmedetomidine for acute post-operative pain following bunionectomy that failed due to lack of observed efficacy. Sponsored by Baudax Bio/Lotus Clinical, the trial was subsequently redesigned and successfully executed as NCT02284243. Table 17 compares the real-world redesign with ClinicalReTrialAgent’s proposals.
Table 17: Real-world validation (NCT02169336, Efficacy Inadequacy): We compare the real-world changes with ClinicalReTrialAgent’s proposed modifications, categorizing alignment as: ✓(perfect match), $\sim$ (strategic alignment, tactical differences), or $\times$ (missed or incorrect).
Change Type Real-World Redesign ClinicalReTrialAgent Proposal Match Impact Level
Major Redesigns (Critical to Efficacy Success)
Statistical Power INCREASED sample size: 95 $\rightarrow$ 168 participants (+77%) INCREASE to $\sim$100 participants (power_multiplier=1.0x) $\sim$ Correct direction, underestimated magnitude
Minor Refinements (Non-Critical to Success)
Primary Outcome KEPT SPID48 unchanged KEEP SPID48 as primary outcome Preserved endpoint
Dosing Regimen KEPT identical (35mcg & 50mcg q6h) KEEP existing 35/50mcg dosing No modifications
Enrichment Criteria KEPT (no biomarker screening) ADD Central Sensitization Inventory (CSI $\geq$50) on top of existing criteria $\times$ Unnecessary restrictiveness (would exclude 80-85%)
Enrichment Criteria KEPT (no biomarker screening) ADD BDNF levels ($\geq$15 ng/ml) on top of existing criteria $\times$ Over-engineered (would exclude 80%)
Enrichment Criteria KEPT (no genetic screening) ADD COMT Val158Met polymorphism screening on top of existing criteria $\times$ Invalid (flagged by validation, would exclude 70%)
The primary cause of trial failure was insufficient statistical power to detect the treatment effect, with only 95 participants enrolled. Both the real-world redesign and ClinicalReTrialAgent correctly identified underpowering as the root cause and proposed sample size increase as the primary solution, demonstrating strong diagnostic capability. However, the real-world approach implemented a single, decisive change—increasing enrollment to 168 participants (+77%)—while maintaining 100% protocol fidelity across eligibility criteria, primary outcomes, and dosing. In contrast, ClinicalReTrialAgent underestimated the required sample size (proposing $\sim$100 vs. actual 168, representing only a 5% increase) and additionally proposed layering biomarker enrichment criteria atop the existing protocol. The agent simultaneously proposed adding three unecessary new enrichment requirements. Notably, the agent’s own validation system flagged the COMT polymorphism proposal as invalid due to insufficient evidence. This case illustrates a critical limitation: while ClinicalReTrialAgent exhibits strong strategic reasoning (correct root cause identification, appropriate prioritization of power), it defaults to mechanistic over-optimization when pragmatic simplicity proves more effective. The real-world success through power-only expansion—requiring zero design complexity—validates Occam’s Razor in trial redesign: sometimes “more participants” decisively outperforms “smarter selection.”
Implications.
The case study reveals that ClinicalReTrialAgent excels at strategic-level redesign (identifying root causes, removing barriers, preserving safety constraints) but lacks tactical-level domain expertise (selecting specific biomarkers, anticipating data quality needs, distinguishing between validity-preserving and validity-threatening modifications). This suggests that future work should integrate specialized biomarker databases and safety constraint ontologies to bridge the gap between strategic reasoning and actionable clinical knowledge.

### v2

@@
-ClinicalReTrial: A Self-Evolving AI Agent for Clinical Trial Protocol Optimization
+ClinicalReTrial: Clinical Trial Redesign with Self-Evolving Agents
 Abstract
-Clinical trial failure remains a central bottleneck in drug development, where minor protocol design flaws can irreversibly compromise outcomes despite promising therapeutics.
-Although cutting-edge AI methods achieve strong performance in predicting trial success, they are inherently reactive for merely diagnosing risk without offering actionable remedies once failure is anticipated.
-To fill this gap, this paper proposes ClinicalReTrial, a self-evolving AI agent framework that addresses this gap by casting clinical trial reasoning as an iterative protocol redesign problem.
-Our method integrates failure diagnosis, safety-aware modification, and candidate evaluation in a closed-loop, reward-driven optimization framework.
-Serving the outcome prediction model as a simulation environment, ClinicalReTrial  enables low-cost evaluation of protocol modifications and provides dense reward signals for continuous self-improvement.
-To support efficient exploration, the framework maintains hierarchical memory that captures iteration-level feedback within trials and distills transferable redesign patterns across trials.
-Empirically, ClinicalReTrial  improves $83.3\%$ of trial protocols with a mean success probability gain of $5.7\%$, and retrospective case studies demonstrate strong alignment between the discovered redesign strategies and real-world clinical trial modifications.
-ClinicalReTrial: A Self-Evolving AI Agent for Clinical Trial Protocol Optimization
-Sixue Xing1, Xuanye Xia2, Kerui Wu3, Meng Jiang1, Jintai Chen4, Tianfan Fu5 1University of Notre Dame, South Bend, IN, USA 2Georgia Institute of Technology, Atlanta, GA, USA 3University of Massachusetts Amherst, Amherst, MA, USA 4Hong Kong University of Science and Technology (Guangzhou), Guangzhou, Guangdong, China 5Nanjing University, Nanjing, Jiangsu, China Correspondence: futianfan@gmail.com
+Clinical trials constitute a critical yet exceptionally challenging and costly stage of drug development (\$2.6B per drug), where protocols are encoded as complex natural language documents, motivating the use of AI systems beyond manual analysis.
+Existing AI methods accurately predict trial failure, but do not provide actionable remedies.
+To fill this gap, this paper proposes ClinicalReTrial, a multi-agent system that formulates clinical trial optimization as an iterative redesign problem on textural protocols.
+Our method integrates failure diagnosis, safety-aware modifications, and candidate evaluation in a closed-loop, reward-driven optimization framework.
+Serving the outcome prediction model as a simulation environment, ClinicalReTrial enables low-cost evaluation and dense reward signals for continuous self-improvement.
+We further propose a hierarchical memory that captures iteration-level feedback within trials and distills transferable redesign patterns across trials.
+Empirically, ClinicalReTrial improves $83.3\%$ of trial protocols with a mean success probability gain of $5.7\%$ with negligible cost (\$0.12 per trial).
+Retrospective case studies demonstrate alignment between the discovered redesign strategies and real-world clinical trial modifications.
+The code is anonymously available at: https://github.com/xingsixue123/ClinicalFailureReasonReTrial.
+ClinicalReTrial: Clinical Trial Redesign with Self-Evolving Agents
+Sixue Xing1, Kerui Wu2, Xuanye Xia3, Meng Jiang1, Jintai Chen4, Tianfan Fu5 1University of Notre Dame, South Bend, IN, USA 2University of Massachusetts Amherst, Amherst, MA, USA 3Georgia Institute of Technology, Atlanta, GA, USA 4Hong Kong University of Science and Technology (Guangzhou), Guangzhou, Guangdong, China 5Nanjing University, Nanjing, Jiangsu, China Correspondence: futianfan@gmail.com
 1 Introduction
-Clinical trials represent the most critical and expensive phase in drug discovery, with an estimated cost of \$2.6 billion (dimasi2016innovation) per approved drug, and low success rates of approximately 10-20% (yamaguchi2021approval).
-Serving as documented plan that specifies the study’s objectives, clinical trial protocols involve complex, interdependent design choices (getz2017trends), such as eligibility criteria, dosing strategies, and endpoint definitions, where small design flaws can propagate into irreversible failure.
-These challenges motivate the use of AI systems zhang2023harnessing that can reason over high-dimensional trial designs, leverage historical evidence, and systematically assess failure risks at scale.
+Clinical trials represent the most critical and expensive phase in drug discovery, with an estimated cost of \$2.6 billion (DiMasi et al., 2016) per approved drug, and low success rates of approximately 10-20% (Yamaguchi et al., 2021).
+Serving as documented plan that specifies the study’s objectives, clinical trial protocols involve complex, interdependent design choices (Getz and Campo, 2017) expressed in natural language documents, such as eligibility criteria, dosing strategies, and endpoint definitions, where small design flaws can propagate into irreversible failure.
+These challenges motivate the use of AI systems Zhang et al.
+(2023) that can reason over high-dimensional trial designs, leverage historical evidence, and systematically assess failure risks at scale.
 Recent advances in AI have enabled increasingly accurate prediction of clinical trial outcomes.
-For example, lo2019machine uses structured metadata to model success likelihood; fu2022hint; chen2024uncertainty; chen2025trialbench integrate heterogeneous data sources using architectures including graph neural networks and hierarchical attention mechanisms to achieve strong predictive performance; yue2024clinicalagent; liu2025autoct incorporate Large Language Models (LLMs) and external knowledge bases to enhance reasoning and explainability in trial outcome prediction.
+For example, Lo et al.
+(2019) uses structured metadata to model success likelihood; Fu et al.
+(2022); Chen et al.
+(2024c, 2025) integrate heterogeneous data sources using architectures including graph neural networks and hierarchical attention mechanisms to predict trial approval rate; Yue et al.
+(2024); Liu et al.
+(2025) incorporate Large Language Models (LLMs) and external knowledge bases to enhance reasoning and explainability in trial outcome prediction through natural advanced language understandings of complex medical texts.
 Despite their success, existing approaches are inherently reactive in nature: they operate on a fixed clinical trial protocol and produce a prediction or post-hoc explanation of trial success or failure.
 However, these methods do not address a more practically consequential problem: they are unable to respond to a determined trial failure due to the lack of actionable interventions.
-In real-world drug discovery, stakeholders require not only assessments of failure risk, but also actionable guidance on protocol redesign, including principled modifications or augmentations informed by the identified sources of risk.
-To bridge this gap, in this work, we propose ClinicalReTrial, a self-evolving AI agent framework that moves beyond static prediction toward actionable intervention via end-to-end trial protocol optimization, while continuously improves its redesign policies.
-The framework instantiates a coordinated multi-agent pipeline that performs failure diagnosis, protocol redesign, and candidate evaluation, with domain knowledge and safety awareness embedded at each decision stage.
+In real-world drug discovery, stakeholders require not only assessments of failure risk, but also actionable guidance on protocol redesign, including principled modifications or augmentations informed by the identified sources of risk (Dagenais et al., 2022; Baumfeld Andre et al., 2020).
+Figure 1: ClinicalReTrial Agent architecture.
+The system operates through iterative refinement: agents analyze failures, generate modifications, and receive rewards from the simulation environment.
+Historical explorations are extracted into structured knowledge that guides subsequent iterations, enabling progressive improvement.
+To bridge this gap, in this work, as shown in Figure˜1, we propose ClinicalReTrial, a self-evolving AI agent that moves beyond static prediction toward actionable intervention via end-to-end trial protocol optimization, while continuously improving its redesign policies.
+The inherently language-rich nature of clinical trial protocols makes this an ideal domain for LLM-based optimization.
+Our framework instantiates a coordinated multi-agent pipeline that performs failure diagnosis, protocol redesign, and candidate evaluation, with domain knowledge and safety awareness embedded at each decision stage.
 Beyond a single optimization run, we adopt the prediction model as a simulation environment to provide rewards for continuous self-improvement.
-Specifically, ClinicalReTrial  maintains local memory to accumulate iteration-level feedback and reward-attributed modification outcomes for within-trial adaptation, while a global memory distills transferable redesign patterns across trials to enable warm start initialization and exploration calibration.
+Specifically, ClinicalReTrial maintains local memory to accumulate iteration-level feedback and reward attributed modification outcomes for within-trial adaptation, while a global memory distills transferable redesign patterns across trials to enable warm start initialization and exploration calibration.
 Through this hierarchical learning structure and reward-driven closed-loop optimization, ClinicalReTrial systematically explores the protocol modification space and learns to identify high-impact interventions that improve clinical trial success probability.
 Experimentally, our prediction model demonstrated the strongest performance (PR-AUC>0.75>0.75), allowing it to serve as a reliable simulation environment for evaluation and agent optimization.
-In the trial redesign experiments, ClinicalReTrial successfully improve $89.3\%$ of trial protocols with mean probability gain $\Delta p=5.7\%$, achieved at negligible cost ($\mathdollar 0.12/trial$).
+In the trial redesign experiments, ClinicalReTrial successfully improved $83.3\%$ of trial protocols with mean probability gain $\Delta p=5.7\%$, achieved at negligible cost ($\mathdollar 0.12/$trial).
 We further conduct multiple real-world retrospective case studies.
-Impressively, the redesigns generated by ClinicalReTrial  exhibit strong strategic alignment with independently derived real-world trial modifications, highlighting the potential of self-evolving AI agents to support principled, clinically grounded trial redesign.
-Main contributions are listed as follows: (1) (to the best of our knowledge) We are the first to formulate clinical trial optimization as an AI-solvable and in silico-verifiable problem.
-(2) We propose a multi-agent pipeline with domain knowledge that decomposes clinical trial protocol optimization into analysis, augmentation, and evaluation.
-(3) We develop a simulation-driven clinical trial optimization framework with hierarchical memory utilization for continuous self-improvement.
+Impressively, the redesigns generated by ClinicalReTrial exhibit strategic alignment with independently derived real-world trial modifications, highlighting the potential of self-evolving AI agents to support principled trial redesign.
+Main contributions.
+are listed as follows: (1) (to the best of our knowledge) We are the first to formulate clinical trial optimization as an AI-solvable and in silico-verifiable problem.
+(2) We propose a multi-agent pipeline with domain knowledge that decomposes clinical trial protocol optimization into diagnosis, modification, and evaluation.
+(3) We develop a simulation-driven clinical trial optimization framework with In-Context Learning and multi-level memory for continuous self-improvement through reward attributed prompt optimization, dynamic redesign pool curation, and cross-trial knowledge distillation.
 2 Related Work
-Early efforts employed classical machine learning (logistic regression (LR), random forests) on expert-curated features (gayvert2016clinical; lo2019machine), establishing feasibility but lacking multi-modal data integration.
-Deep learning approaches addressed this: fu2022hint proposed HINT, integrating drug molecules, ICD-10 codes, and eligibility criteria; chen2024uncertainty added uncertainty quantification; wang2024twin designed LLM-based patient-level digital twins; chen2025trialbench released a standardized TrialBench with multi-modal baselines.
-while maintaining competitive performance.
-Recent LLM approaches demonstrate medical reasoning (singhal2023medpalm), enhanced via retrieval-augmented generation (lewis2020rag) with databases like DrugBank (wishart2018drugbank), Hetionet (himmelstein2017hetionet), and domain-adapted encoders like BioBERT (lee2020biobert).
-Building on this, yue2024clinicalagent introduced ClinicalAgent, decomposing prediction into specialized sub-task agents with ReAct reasoning (yao2023react).
-liu2025autoct proposed AutoCT for autonomous feature engineering via Monte Carlo Tree Search (chi2024sela).
-However, these methods function as discriminators mapping protocols to success probabilities without explaining why failures occur or how to modify protocols.
-First to formulate generative optimization, our multi-agent architecture leverages chain-of-thought (wei2022chain), and least-to-most prompting (zhou2023least) for hierarchical problem decomposition.
+Early efforts employed classical machine learning (logistic regression (LR), random forests) on expert-curated features (Gayvert et al., 2016; Lo et al., 2019), establishing feasibility but lacking multi-modal data integration.
+Deep learning approaches addressed this: Fu et al.
+(2022) proposed HINT, integrating drug molecules, ICD-10 codes, and eligibility criteria; Chen et al.
+(2024c) added uncertainty quantification and interpretability; Wang et al.
+(2024) designed LLM-based patient-level digital twins; Chen et al.
+(2025) released a standardized TrialBench with multi-modal baselines while maintaining competitive performance.
+Recent LLM approaches demonstrate medical reasoning (Singhal et al., 2023), enhanced via retrieval-augmented generation (Lewis et al., 2020) with databases like DrugBank (Wishart et al., 2018), Hetionet (Himmelstein et al., 2017), and domain-adapted encoders like BioBERT (Lee et al., 2020).
+Building on this, Yue et al.
+(2024) introduced ClinicalAgent, decomposing prediction into specialized sub-task agents with ReAct reasoning (Yao et al., 2023).
+Liu et al.
+(2025) proposed AutoCT for autonomous feature engineering via Monte Carlo Tree Search (Chi et al., 2024).
+However, these methods are essentially predictive models without explaining why failures occur or how to modify protocols.
+In contrast, our multi-agent architecture leverages chain-of-thought (Wei et al., 2022) and least-to-most prompting (Zhou et al., 2023) to dynamically diagnose failure causes and refine protocols.
 3 Methodology
-Overview
-We propose ClinicalReTrial, a self-improving multi-agent system that redesigns failed clinical trials through reward-driven iterative optimization.
-First, Section 3.1 formulates the clinical trial optimization problem.
-Then, Section 3.2 details the agent components, Section 3.3 presents the hierarchical learning mechanisms, and Section 3.4 describes the knowledge retrieval system.
-For ease of exposition, Figure 1 illustrates the whole process and Algorithm 1 formalizes the iterative optimization procedure.
-Figure 1: ClinicalReTrial Agent architecture.
-The system operates through iterative refinement: agents analyze failures, generate augmentations, and receive rewards from the simulation environment.
-Modifications are distilled into structured knowledge that guides subsequent iterations, enabling progressive improvement.
-Algorithm 1 Clinical trial protocol optimization with hierarchical learning.
-0: Failed trial $T_{0}$, failure mode $y$, global memory $\mathcal{M}^{\text{global}}$
+Overview.
+ClinicalReTrial is a self-improving multi-agent system that redesigns failed clinical trials through reward-driven iterative optimization.
+Specifically, we first formulate the clinical trial optimization problem in Section˜3.1.
+Then, we build a multi-agent framework to address it in Section˜3.2.
+To further improve our Agent performance, we integrate a knowledge retrieval system and present In-Context Learning with multi-level memory in Section˜3.3.
+For ease of exposition, Figure˜2 illustrates the whole process and Algorithm˜1 provided in Appendix A formalizes the iterative optimization procedure.
+Figure 2: Detailed ClinicalReTrial Agent architecture and iterative redesign workflow, comprising diagnosis, augmentation, and validation, with a simulation environment and hierarchical memory to progressively optimize failed clinical trial protocols.
+3.1 Problem: Clinical Trial Optimization
+The goal of clinical trial is to evaluate the safety and efficacy of drug on patients.
+A clinical trial protocol, including eligibility criteria and drug dosage specifications, defines the study’s framework in a standardized and controlled manner.
+Formally, let $T_{0}=\{e_{1},e_{2},\ldots,e_{K}\}$ denote a clinical trial protocol decomposed into $K$ modifiable elements (e.g., eligibility criteria, dosage regimens, endpoint definitions), where a prediction model $f_{\theta}:\mathcal{T}\rightarrow[0,1]$ assigns success probability $p_{0}=f_{\theta}(T_{0})$.
+Each element $e_{i}$, after redesign, admits a set of augmentations $\mathcal{A}_{i}=\{a_{i1},a_{i2},\ldots,a_{im_{i}}\}$ representing clinically valid modifications, where candidate protocols $\mathbb{T}=\{T^{\prime}_{1},T^{\prime}_{2},\ldots,T^{\prime}_{N}\}$ are reconstructed by selectively replacing elements in $T_{0}$ with their augmented variants.
+Such an exploration set of redesigns $\mathbb{T}$ is evaluated through the prediction model to obtain success probabilities, with each $p^{\prime}_{j}=f_{\theta}(T^{\prime}_{j})$.
+The optimization objective seeks the optimal protocol $T^{*}=\arg\max_{T^{\prime}\in\mathbb{T}}f_{\theta}(T^{\prime})$ with success probability $p^{*}$, where overall improvement is measured by $\Delta p=p^{*}-p_{0}$ with $p^{*}=f_{\theta}(T^{*})$.
+3.2 Multi-Agent Architecture
+With heterogeneous necessities, clinical trial redesign requires distinct diagnostic expertise and remediation strategies Fogel (2018).
+We introduce a multi-agent architecture Chen et al.
+(2024b) that naturally align with this structure, comprising four coordinated components: Trial Failure Analyzer (§3.2.1) performs root cause diagnosis; the Protocol Refinement Generator (§3.2.2) synthesizes modifications; the Clinical Safety Validator (§3.2.3) prunes unsafe modifications; and the Protocol Candidate Evaluator (§3.2.4) provides simulation based feedback.
+3.2.1 Protocol Diagnosis Analyzer
+Given a failed protocol and prior failure modes: {Poor Enrollment, Safety/Adverse Effect, Drug Lack of Efficacy}, the Analyzer Agent produces a prioritized set of modifications based on Protocol Taxonomy and Action Determination, targeting the protocol feature under consideration; then specifies the action strategy {Delete, Modify, Add} with confidence score.
+Protocol Taxonomy.
+Before being processed through the agent redesign pipeline, protocol features are classified using medical terminology into structured categories: eligibility criteria, safety exclusions, selection criteria, and enrichment criteria.
+We further categorized dosage and outcomes by safety risk, failure contribution, and modification efficacy.
+These classifications guide action selection based on the observed failure reason and its corresponding analysis.
+Action Determination.
+The agent extracts failure signatures via action category alignment and confidence scoring, calibrated using historical modification success patterns.
+At iteration $t=1$, the agent receives warm start guidance from cross-trial memory (§3.3.3); from $t\geq 2$, it incorporates performance patterns from prior iterations (§3.3.2).
+These insights yield prioritized modification targets balancing domain knowledge with empirical feedback, which guide the Generator Agent (§3.2.2) to produce concrete modifications.
+3.2.2 Protocol Refinement Generator
+The Generator Agent translates diagnostic insights from the Analyzer Agent into diverse design refinements that address identified weaknesses while preserving clinical validity.
+Action-specific Variant Generation.
+The agent employs action-specific logic: Delete critical failure factors while preserving safety; Modify adjusts thresholds or operationalizes vague terms; Add introduces biomarker enrichment or contraindication criteria.
+Various modifications are first validated for clinical safety (§3.2.3), then proceed to formal evaluation.
+3.2.3 Clinical Safety Validator
+To ensure the proposed augmentations satisfy clinical safety standards, the system uses LLM-as-a-Judge (Zheng et al., 2023) with domain database retrieval and autonomous validation.
+Database Retrieval.
+The multi-agent system enhances embedded parametric knowledge with targeted retrieval from biomedical databases: DrugBank (Wishart et al., 2018) with pharmacological profiles including toxicity, metabolism, contraindications; Disease Database (Chen et al., 2024a) that contains diagnostic criteria, symptomatology, risk factors; and PubMed Abstract spanning 1975-2025.
+Retrieval employs dense embeddings (BioBERT (Lee et al., 2020) for drugs/diseases, PubMedBERT (Gu et al., 2021) for literature) with FAISS indexing.
+Retrieved results have tangential content filtered, while enforcing strict temporal constraints that limit PubMed queries to prevent outcome leakage.
+Autonomous Safety Validation.
+The agent checks and prunes unsafe modifications (dosage changes, population shifts, contraindications) triggering retrieval from external databases when identifying knowledge gaps in medical assessments.
+Validated candidates that passed the validation proceed to the Evaluator (§3.2.4) for simulation-based evaluation.
+3.2.4 Protocol Candidate Evaluator
+The Evaluator combines validated augmentations with the original trial protocol into complete redesigned trial candidates, each evaluated through simulation-based assessment with a success probability assigned that guides modification and hierarchical learning (§3.3).
+Search Strategy.
+Candidate trials are formed by combining validated augmentations across original protocols, where Beam Search Freitag and Al-Onaizan (2017) is used to reduce exponential complexity to approximately quadratic.
+Simulation Environment.
+To provide reliable feedback for agent-generated modifications, we train a model that predicts trial candidates’ outcome probabilities from encoded trial features, serving as a simulation environment that enables rapid evaluation for thousands of redesigns without conducting actual clinical trials, guiding the agent system toward promising protocol optimization.
+The improvement in predicted success probability, from the original trial to the redesigned candidate, serves as the reward signal.
+3.3 In-Context Learning with Multi-level Memory
+Without memory or iterative feedback, multi-agent systems repeat failed strategies and cannot leverage historical performance.
+We address this via In-Context Learning Brown et al.
+(2020) from reward guided feedback (§3.3.1) with knowledge consolidation operating at two temporal scales: within-trial learning3.3.2) accumulates local memory across iterations for trial refinement, while cross-trial learning3.3.3) maintains global memory to transfer successful patterns across the trial corpus.
+3.3.1 Redesign Reward
+To identify specific modifications that drive improvement, we decompose protocol-level redesigns outcome probabilities $p(T^{\prime})$ into augmentation-level rewards.
+The Evaluator first evaluates combined trial variants via prediction, then attributes credit to individual modifications.
+For each augmentation $m$, we compute its marginal contribution $r(m)$ across the explored combinatorial space:
+$r(m)=\mathbb{E}_{m\in T^{\prime}}[p(T^{\prime})]-\mathbb{E}_{m\not\in T^{\prime}}[p(T^{\prime})].$ (1)
+The complete reward distribution $\mathcal{R}_{t}$ encompasses all redesign augmentations with their validation status and rewards, enabling performance stratified knowledge extraction from both successful modifications and contraindicated patterns.
+3.3.2 Within-trial Learning: Local Iterative Optimization
+To refine LLM behavior by prompt optimization Ramnath et al. (2025), we extract short-term memory from reward $\mathcal{R}_{t}$ and integrate with Agent.
+Knowledge Extraction.
+We partition $\mathcal{R}_{t}$ into two types of knowledge: action-level patterns $\mathcal{K}^{s}_{t}$ aggregate modification type performance across aspects; and example-level demonstrations $\mathcal{K}^{t}_{t}$ comprise performance stratified modifications.
+Agent Integration.
+The Analyzer’s aspect prioritization logic integrates $\mathcal{K}^{s}_{t}$ via coverage-based confidence scoring: penalizing repeated patterns, rewarding unexplored spaces, and weighting by historical success rates.
+The Generator Agent samples performance stratified exemplars $\mathcal{K}^{t}_{t}$ for few-shot prompting Brown et al.
+(2020).
+While the Evaluator maintains a redesign pool of high performing and positively rewarded modifications for combinatorial search reuse, enabling test-time search space scaling Snell et al.
+(2024).
+3.3.3 Cross-trial Learning: Global Memory
+After each trial converges, global memory is extracted via LLM synthesis, where each trial benefits from and contributes to the evolving knowledge pool.
+Building on meta-learning frameworks Parisi et al.
+(2019), we implement cross-trial knowledge transfer, maintaining generalizable patterns via two representations: (1) Qualitative Strategic Guidance.
+Aspect-level recommendations extracted from high-performing redesign patterns provide warm start initialization for the Analyzer Agent.
+(2) Quantitative Statistical Signatures.
+Recorded mean reward, variance, and modification success rates enable the Generator Agent to calibrate exploration intensity, scaling generation count inversely with historical success rates and proportionally to pattern variance.
+4 Experiment
+We evaluate ClinicalReTrial across two dimensions: (1) simulation environment performance, validating that simulation environment achieve sufficient accuracy to serve as reliable feedback oracles, and (2) Agent optimization quality, demonstrating that our multi-agent system successfully redesigns failed trials through iterative learning.
+4.1 Experimental Setup
+Our system is built on GPT-4o-mini and evaluated on failed clinical trials from the TrialBench dataset (Chen et al., 2025).
+Using 20769 annotated Phase I-IV trials, we encode multi-modal features into 6,173-dimensional embeddings (details in Appendix C.2) and train LightGBM (Ke et al., 2017) classifiers to predict trial outcome $\hat{y}\in[0,1]$.
+We follow TrialBench’s train-test split, further splitting the training set 8:2 for training-validation.
+Due to computational constraints, we evaluate the agent on random selected sample of 60 failed trials from the test set (20 enrollment, 20 safety, 20 efficacy failures) representing diverse trial phases (Data detail in Appendix C.1).
+The agent operates with a 5-iteration budget.
+We use exhaustive search when the combinatorial space has <1,000<1{,}000 candidates, otherwise beam search with width $k=8$.
+We measure effectiveness through predicted probability improvement, threshold achievement rate, and convergence efficiency.
+The code is anonymously available at: https://github.com/xingsixue123/ClinicalFailureReasonReTrial.
+Table 1: Performance comparison across three binary prediction tasks.
+Task Model ROC-AUC ($\uparrow$) PR-AUC ($\uparrow$) Fail Detection ($\uparrow$) Poor Enrollment TrialBench 0.613 $\pm$ 0.007 0.626 $\pm$ 0.011 0.525 $\pm$ 0.013 HINT 0.534 $\pm$ 0.010 0.613 $\pm$ 0.012 0.580 $\pm$ 0.018 Logistic Reg.
+0.622 $\pm$ 0.010 0.696 $\pm$ 0.012 0.669 $\pm$ 0.012 ClinicalReTrial 0.676 $\pm$ 0.009 0.754 $\pm$ 0.010 0.740 $\pm$ 0.012 Drug Adverse Effect TrialBench 0.587 $\pm$ 0.017 0.892 $\pm$ 0.006 0.427 $\pm$ 0.035 HINT 0.513 $\pm$ 0.014 0.882 $\pm$ 0.009 0.459 $\pm$ 0.031 Logistic Reg.
+0.612 $\pm$ 0.018 0.909 $\pm$ 0.008 0.422 $\pm$ 0.029 ClinicalReTrial 0.656 $\pm$ 0.018 0.925 $\pm$ 0.007 0.695 $\pm$ 0.028 Drug Efficacy TrialBench 0.692 $\pm$ 0.012 0.862 $\pm$ 0.006 0.565 $\pm$ 0.020 HINT 0.559 $\pm$ 0.013 0.841 $\pm$ 0.008 0.525 $\pm$ 0.021 Logistic Reg.
+0.665 $\pm$ 0.015 0.886 $\pm$ 0.009 0.549 $\pm$ 0.025 ClinicalReTrial 0.746 $\pm$ 0.013 0.914 $\pm$ 0.007 0.725 $\pm$ 0.021
+Table 2: Performance of multi-class clinical trial outcome prediction across trial phases.
+Model Phase 1 Phase 2 Phase 3 Phase 4 ROC-AUC ($\uparrow$) PR-AUC ($\uparrow$) ROC-AUC ($\uparrow$) PR-AUC ($\uparrow$) ROC-AUC ($\uparrow$) PR-AUC ($\uparrow$) ROC-AUC ($\uparrow$) PR-AUC ($\uparrow$) TrialBench 0.475 $\pm$ 0.027 0.255 $\pm$ 0.006 0.569 $\pm$ 0.010 0.295 $\pm$ 0.008 0.550 $\pm$ 0.012 0.279 $\pm$ 0.008 0.477 $\pm$ 0.022 0.256 $\pm$ 0.007 HINT 0.540 $\pm$ 0.022 0.272 $\pm$ 0.009 0.535 $\pm$ 0.009 0.267 $\pm$ 0.005 0.474 $\pm$ 0.019 0.251 $\pm$ 0.006 0.548 $\pm$ 0.021 0.273 $\pm$ 0.014 Logistic Reg.
+0.606 $\pm$ 0.019 0.326 $\pm$ 0.015 0.583 $\pm$ 0.011 0.306 $\pm$ 0.008 0.621 $\pm$ 0.016 0.350 $\pm$ 0.017 0.550 $\pm$ 0.022 0.280 $\pm$ 0.010 ClinicalReTrial 0.633 $\pm$ 0.016 0.344 $\pm$ 0.016 0.662 $\pm$ 0.011 0.382 $\pm$ 0.011 0.669 $\pm$ 0.017 0.412 $\pm$ 0.019 0.543 $\pm$ 0.025 0.282 $\pm$ 0.012
+4.2 Simulation Environment Performance
+Our simulation environment model is compared against baseline approaches, including TrialBench (Chen et al., 2025) and HINT (Fu et al., 2022), prior state-of-the-art systems operating on the original TrialBench feature space.
+Additionally, a Logistic Regression model is trained on the same encoded features (Appendix C.2) for reference.
+Failure-specific Prediction.
+Implemented in ClinicalReTrial Agent, the simulation environment must correctly predict specific failure outcomes.
+We train three independent models on our encoded features, each targeting one failure detection task against success.
+Table˜1 report comprehensive metrics across our models and baseline approaches trained for the task: Poor Enrollment, Safety/Adverse Effect and Lack of Efficacy prediction.
+Our model achieves PR-AUC >0.75>0.75 across all failure modes, providing reliable discriminative feedback.
+All models achieve Failure Detection Rates of 70-74% with task-specific prediction thresholds ($p\geq 0.6$ for enrollment, $p\geq 0.9$ for safety, $p\geq 0.85$ for efficacy).
+This ensures that predicted probability shifts $\Delta p>0.03$ indicate improved trial designs (Appendix C.1).
+TrialBench Benchmark.
+We further validate the same model architecture against existing benchmarks on the TrialBench 4-class classification task (predicting Success, Enrollment Failure, Safety Failure, or Efficacy Failure) (Chen et al., 2025).
+Table˜2 shows that the base model used in our simulation environment outperformed all baselines, achieving a ROC-AUC improvement of $6\%$ to $19\%$.
+Feature Importance Analysis.
+We studied feature importance analysis with SHAP (Lundberg and Lee, 2017).
+Consistent with our hypothesis, eligibility, drug-disease interaction features, and endpoint alignment features are most important for outcomes prediction (Appendix C.3).
+4.3 Protocol Optimization - Quantitative Results
+Having confirmed the simulation environment’s reliability, we evaluate ClinicalReTrial Agent’s ability to redesign failed clinical trials.
+Convergence Analysis.
+Table˜3 reports comprehensive convergence statistics across all trials.
+The Agent improved 83.3% of protocol designs (50/60 successfully processed trials showed positive $\Delta p$), with 4 trials (6.7%) encountering agent failures where the system identified zero opportunities of potential redesign, occurring in the efficacy failure mode.
+Table 3: Convergence and probability shift analysis by failure mode.
+Failures denotes trials where the agent identified no redesign opportunities; Threshold reports trials exceeding the predefined improvement threshold.
+$p_{0}$$p_{\text{final}}$: initial/optimized success probabilities; IQR: 25th–75th percentile of $\Delta p$.
+Mode Trials Failures Improved Trials Threshold $p_{0}$ / $p_{\text{final}}$ $\Delta p$ IQR Enrollment 20 0 20/20 (100%) 8/20 (40%) 0.506 / 0.563 +0.058 [+0.034, +0.068] Safety 20 0 18/20 (90%) 4/20 (20%) 0.791 / 0.831 +0.070 [+0.032, +0.092] Efficacy 20 4 12/20 (60%) 10/20 (50%) 0.813 / 0.859 +0.039 [+0.013, +0.040] Overall 60 4 (6.7%) 50/60 (83.3%) 22/60 (36.7%) 0.695 / 0.730 +0.057 [+0.029, +0.073]
+Table 4: Ablation study of the ClinicalReTrial framework.
+Each component is individually replaced with a plain LLM call to isolate its contribution, together with a CoT baseline.
+Stats computed using paired $t$-tests with $n=10$ enrollment failure trials.
+*$p<0.05$, **$p<0.01$, ns: not significant.
+Iter5 dashes indicate baselines that lack iterative capability and are evaluated at iteration 1 only.
+Method N Iter1 $\Delta p$ Iter5 $\Delta p$ $p$-value Cohen’s $d_{z}$ Test Basis Full System 10 $037\pm 0.024 +0.0$$057\pm 0.023$ w/o Memory 10 $024\pm 0.018 +0.0$$38\pm 0.016 0.0$$007^{**}$ $1.10$ Iter5 vs Full w/o Pool 10 $037\pm 0.024 +0.0$$45\pm 0.022 0.0$$005^{**}$ $1.18$ Iter5 vs Full w/o Augmentor 10 $025\pm 0.032 +0.0$$050\pm 0.028$ $0.189^{\text{ns}}$ $0.45$ Iter5 vs Full w/o Planner 10 $026\pm 0.017 0.0$$048^{*}$ $0.94$ Iter1 vs Full CoT 10 $008\pm 0.009 0.0$$009^{**}$ $1.14$ Iter1 vs Full
+The system demonstrated efficient convergence patterns, with 15% (9/60) of trials exhibiting natural termination before iteration 5 due to exhausted modification space.
+Most trials used all 5 iterations, suggesting adaptive stopping could improve efficiency.
+Iterative Improvement Trajectories.
+We examine the learning trajectory of successfully processed trials (56/60).
+The system achieved a mean improvement of $\Delta p=+0.057$.
+Table˜3 stratifies results by failure mode, while Figure˜3 illustrates learning dynamics by iteration.
+Performance heterogeneity across failure modes reflects the differential amenability of clinical trial design elements to protocol-level intervention.
+Safety failures exhibit the largest improvements (mean $\Delta p=+0.070$, IQR [+0.032, +0.092]), as adverse events often stem from identifiable contraindication patterns that can be systematically addressed through eligibility refinement and dosage adjustment.
+Enrollment failures show substantial gains (mean $\Delta p=+0.058$).
+Efficacy failures demonstrate the smallest yet statistically meaningful improvements (mean $\Delta p=+0.040$), as therapeutic effectiveness depends heavily on drug-disease compatibility, where sometimes protocol modifications alone cannot compensate for fundamental drug inefficacy.
+Figure 3: Iterative redesign trajectories by failure mode. Curves show mean cumulative improvement in predicted success probability ($\Delta p$).
+As shown in Figure˜3, the learning trajectory reveals major initial gains followed by decreasing returns.
+This diminishing returns pattern validates the knowledge distillation mechanism: high-quality modifications are identified early through rapid retrieval boosted analysis, while later iterations exploit narrower optimization opportunities by refining secondary parameters or addressing edge case contraindications.
+Computational Cost and Efficiency.
+The system demonstrates practical feasibility with a mean cost of \$0.12 per trial across 56 trials, which is a negligible fraction of typical \$2.6 billion drug Linear scaling enables industrial deployment: 1,000 trials cost  \$120, establishing ClinicalReTrial as practical for systematic optimization at scale.
+Full computation cost results are in Appendix B.
+Figure 4: ClinicalReTrial Agent’s flowchart on Poor Enrollment failed trial case study (NCT01298752, 2011-02-16), together with the real-world redesign (NCT01591161, 2012-05-02), demonstrating strategic alignments.
+Ablation Study on Self-improvement.
+To validate architectural contributions, we conducted paired ablation across 10 enrollment failure trials (sufficient to detect large effect sizes, Cohen’s $d_{z}>1.0$, at $\alpha=0.05$ with paired designs), individually replacing each component with a plain LLM call while keeping all other modules intact.
+We also compare against a chain-of-thought (CoT) baseline that bypasses multi-agent decomposition entirely.
+Results are reported in Table˜4.
+Among iterative components, removing memory yields significant degradation at both iteration 1 and 5 (gaps of $0.013$ and $0.019$ in $\Delta p$, respectively), confirming that cross-trial warm start provides immediate benefits that compound over iterations.
+Removing the redesign pool produces comparable degradation at iteration 5 (gap of $0.012$, $p<0.01$), validating the importance of reusing high reward modifications for exploitation.
+Replacing the Augmentor with a plain LLM cause performance degradation but not reach significance ($p=0.189$), suggesting it could be partially compensated by upstream diagnostic guidance and downstream safety filtering.
+4.4 Case Studies - Qualitative Result
+Alignment with Real-World Trial Redesigns.
+We analyze clinical trial pairs, where investigators successfully redesigned and re-executed failed protocols spanning enrollment, safety, and efficacy failure modes against ClinicalReTrial’s redesign to validate its applicability, as well as to provides critical insight into clinical applicability.
+Such trial pair cases are chosen based on the mechanistic interpretability for systematic alignment measures.
+We present a poor enrollment redesign case (safety and efficacy cases in Appendix H): NCT01298752, a Phase-III trial of Mapracorat (anti-inflammatory ophthalmic suspension) for post-cataract surgery inflammation that failed due to slow enrollment.
+Sponsored by Bausch & Lomb, the trial was subsequently redesigned and successfully executed as NCT01591161.
+Figure˜4 illustrates ClinicalReTrial’s iterative refinement process across three optimization cycles.
+The Enrollment barrier (cataract surgery waiting requirement) is efficiently identified with a positive reward provided by the simulation environment.
+The agent also progressively explores the modification space: baseline AC cell requirements are successfully added as an enrichment criterion; while the agent also explores the safety enhancement, but end up failing to fully align with real-world redesign.
+RAG-Grounded Validator Case Studies.
+We further present three validator case studies to highlight the role of RAG in improving the clinical reliability of ClinicalReTrial.
+Across these examples, disease-specific, literature-based, and drug-level retrieval provides external medical evidence that enables the Validator Agent to identify unsafe or misaligned redesigns that would otherwise appear plausible at the language level.
+Full case descriptions are deferred to Appendix D.
+5 Conclusion
+In this work, we present ClinicalReTrial, a novel self-evolving agent that moves beyond passive clinical trial outcome prediction to enable proactive optimization of clinical trial protocols.
+Evaluated on the TrialBench benchmark, ClinicalReTrial achieves strong predictive performance while demonstrating the ability to discover significant protocol improvements with clinical best practices, highlighting the potential of agentic AI systems to serve as practical clinical decision support systems for more efficient autonomous trial design.
+Case studies on trial pairs validate that ClinicalReTrial exhibit strategic alignment with independently derived real-world trial modifications.
+Future work will explore tighter integration with real-world constraints.
+Appendix A ClinicalReTrialAlgorithm
+Algorithm 1 Clinical trial optimization with In-Context Learning and Multi-level Memory.
+0: Failed trial $T_{0}$, failure mode $y\in\{\text{enrollment, safety, efficacy}\}$, global memory $\mathcal{M}^{\text{global}}$
 0: Optimized protocol $T^{*}$, best reward $r_{\text{best}}$
-1: Initialize: $r_{\text{best}}\leftarrow 0$, $T^{*}\leftarrow T_{0}$, $\mathcal{M}_{0}^{\text{local}}\leftarrow\emptyset$
+1: Initialize: $r_{\text{best}}\leftarrow 0$, $T^{*}\leftarrow T_{0}$, $\mathcal{M}_{0}^{\text{local}}\leftarrow\emptyset$, $\mathcal{H}_{0}\leftarrow\emptyset$
 2:for $t=1$ to $N_{\text{max}}$ do
@@
 […]
-6:  if $r_{\text{max}}>r_{\text{best}}$ then $r_{\text{best}}\leftarrow r_{\text{max}}$, $T^{*}\leftarrow\arg\max_{T^{\prime}\in\text{explored}}r(T^{\prime})$
 […]
-8:end for
+7:  if $r_{\text{max}}>r_{\text{best}}$ then $r_{\text{best}}\leftarrow r_{\text{max}}$, $T^{*}\leftarrow\arg\max_{T^{\prime}\in\mathbb{T}}f_{\theta}(T^{\prime})$
 […]
-10:return $T^{*}$, $r_{\text{best}}$
-3.1 Problem: Clinical Trial Optimization
-Let $T_{0}=\{e_{1},e_{2},\ldots,e_{K}\}$ denote a clinical trial protocol decomposed into $K$ modifiable elements (e.g., eligibility criteria, dosage regimens, endpoint definitions), where a prediction model $f_{\theta}:\mathcal{T}\rightarrow[0,1]$ assigns success probability $p_{0}=f_{\theta}(T_{0})$.
-Each element $e_{i}$ admits a set of augmentations $\mathcal{A}_{i}=\{a_{i1},a_{i2},\ldots,a_{im_{i}}\}$ representing clinically valid modifications, constructing candidate protocols $T^{\prime}=T_{0}\oplus\mathcal{S}$ by recombining augmented elements with the original protocol.
-Given the exploration set of redesigned protocol candidates $\mathbb{T}=\{T^{\prime}_{1},T^{\prime}_{2},\ldots,T^{\prime}_{N}\}$ evaluated through the model to obtain predicted probabilities, each $p^{\prime}_{j}=f_{\theta}(T^{\prime}_{j})$, the optimization objective seeks for maximizing success probability ($p^{*}$), with overall improvement measured by $\Delta p=p^{*}-p_{0}$ where $p^{*}=f_{\theta}(T^{*})$.
-3.2 Agent Pipeline
-3.2.1 Analysis: Structured Failure Diagnosis
-The Analysis Agent identifies modification targets within failed trial protocols.
-Given a failed protocol and prior failure modes: {Poor Enrollment, Safety/Adverse Effect, Drug Lack of Efficacy}.
-the agent produces a prioritized set of modification, targeting the protocol feature under consideration.
-Then specifies the action strategy, {Delete, Modify, Add} with confidence score.
-Protocol Taxonomy.
-Protocol features are classified by catagories: eligibility criteria into participation barriers, safety exclusions, selection criteria, and enrichment criteria; dosage/outcomes by safety risk, failure contribution, and modification efficacy.
-This taxonomy guides action selection given the observed failure reason and according analysis.
-Action Determination.
-The agent extracts failure signatures via action category alignment and confidence scoring, pre-calibrated using historical modification success patterns.
-At iteration $t=1$, the agent receives warm start guidance from cross-trial memory (§3.3.3); from $t\geq 2$, it incorporates performance patterns from prior iterations (§3.3.2).
-These insights yield prioritized modification targets balancing domain knowledge with empirical feedback, which guide the Augmentation Agent (§3.2.2) to generate concrete modifications.
-3.2.2 Augmentation: Structurally Adaptive Variant Generation
-The Augmentation Agent translates diagnostic insights from the Analysis Agent into diverse design refinements that address identified weaknesses while preserving clinical validity.
-Action-specific Variant Generation.
-The agent employs action-specific logic: Delete critical failure factors while preserving safety; Modify adjusts thresholds or operationalizes vague terms; Add introduces biomarker enrichment or contraindication criteria.
-Multiple variants per target enable exploration, with candidates proceeding to validation (§3.2.3).
-3.2.3 Self-validation: Progressive Safety Assurance
-To make sure ClinicalReTrialAgent’s proposed augmentations satisfy clinical safety standards, the system employs LLM-as-a-Judge (zheng2023judgingllmasajudgemtbenchchatbot) through hierarchical validation stages.
-Autonomous Safety Validation.
-The agent checks and prunes unsafe modifications (dosage changes, population shifts, contraindications) and autonomously retrieves evidence from DrugBank, Disease Database, and PubMed (§3.4) when parametric knowledge is insufficient.
-Validated candidates that passed all stages of progressive filtering proceed to the Exploration Orchestrator (§3.2.4) for simulation-based evaluation and reward assignment.
-3.2.4 Exploration Orchestrator with Simulation Environment
-The Exploration Orchestrator combines validated augmentations with the original trial protocol into complete redesigned trial candidates, each evaluated through simulation based assessment with outcome probability assigned that guides augmentation rewards and hierarchical learning (§3.3).
-Search Strategy.
-Candidate trials are formed by combining validated augmentations across orginal protocols, where Beam search is used to reduce exponential complexity to approximately quadratic while maintaining candidate quality.
-Simulation Environment.
-To provide reliable feedback for agent generated augmentations, we train model that predicts trial candidates’ outcome probabilities from encoded trial features, serving as simulation environment that enables rapid evaluation for thousands of redesigns without conducting actual clinical trials, guiding the agent system toward promising protocol optimization.
-The improvement in predicted success probability, from the original trial to the redesigned candidate, serves as the reward signal.
-3.3 Hierarchical Learning System
-Our framework enables progressive improvement through hierarchical knowledge consolidation operating at two temporal scales: within-trial learning accumulates local memory $\mathcal{M}_{t}^{\text{local}}$ across iterations for trial-specific refinement, while cross-trial learning maintains global memory $\mathcal{M}^{\text{global}}$ to transfer successful patterns across the trial corpus.
-3.3.1 Redesign Reward
-To identify which individual modifications drive improvement, we decompose protocol-level outcome probabilities into augmentation-level rewards.
-The Exploration Orchestrator first evaluates combined trial variants via prediction, then attributes credit to individual modifications.
-For each augmentation $m$, we compute its marginal contribution across the explored combinatorial space:
-$r(m)=\mathbb{E}_{T^{\prime}\ni m}[p(T^{\prime})]-\mathbb{E}_{T^{\prime}\not\ni m}[p(T^{\prime})].$ (1)
-The complete reward distribution $\mathcal{R}_{t}=\{(m_{i},r(m_{i}),v_{i})\}$ encompasses all augmentations with their rewards and validation status, enabling performance-stratified knowledge extraction from both successful modifications and contraindicated patterns.
-3.3.2 Within-trial Learning: Iterative Optimization
-Knowledge distillation operates on the complete reward distribution $\mathcal{R}_{t}$, serving as short-term memory.
-Strategic-level knowledge $\mathcal{K}^{s}_{t}$ guides which modification categories merit prioritization.
-Tactical modification examples $\mathcal{K}^{t}_{t}$ are used to calibrate how modifications should be formulated.
-Agent Integration.
-The Analysis Agent partitions aspects into three coverage sets: previously failed modifications (zero confidence), previously successful modifications (diversification penalty to avoid over-exploitation), and unexplored modifications (exploration bonus).
-Confidence scores combine coverage-based adjustments with action-type success rates from historical trials.
-The Augmentation Agent samples performance-stratified exemplars for few-shot prompting, while the Exploration Agent maintains a redesign pool of top-quartile modifications ($r>0$, 75th percentile) for combinatorial search reuse.
-3.3.3 Cross-trial Learning: Global Memory
-Global memory $\mathcal{M}^{\text{global}}$ maintains generalizable patterns through two representations: Qualitative strategic guidance (aspect-level recommendations extracted via LLM synthesis from high performing redesign patterns) provides warm start initialization for the Analysis Agent at iteration $t=1$, where iteration feedback is absence; Quantitative statistical signatures (mean reward, variance, modification success rate) enable the Augmentation Agent to continuously calibrate exploration intensity, scaling generation count inversely with historical success rates and proportionally to pattern variance.
-After each trial converges, patterns distilled from the final iteration, enabling systematic knowledge transfer where each trial benefits from and contributes to the evolving global memory.
-3.4 Database Retrieval with Self-reflective Validation
-The multi-agent system augments embedded parametric knowledge with targeted retrieval from curated biomedical databases: DrugBank (wishart2018drugbank) with pharmacological profiles including toxicity, metabolism, contraindications; Disease Database (chen2024cod) that contains diagnostic criteria, symptomatology, risk factors; and PubMed Abstract spanning 1975-2025.
-Retrieval employs dense embeddings (BioBERT (lee2020biobert) for drugs/diseases, PubMedBERT (gu2021domain) for literature) with FAISS indexing.
-Retrieved results undergo LLM-driven validation, filtering tangential content, while enforcing strict temporal constraints that limit PubMed queries to prevent outcome leakage.
-4 Experiment
-We evaluate ClinicalReTrial across two dimensions: (1) simulation environment performance, validating that GBDT-based outcome predictors achieve sufficient accuracy to serve as reliable feedback oracles, and (2) ClinicalReTrialAgent optimization quality, demonstrating that our multi-agent system successfully redesigns failed trials through iterative learning.
-4.1 Experimental Setup
-Our system is built on GPT-4o-mini and evaluated on failed clinical trials from the TrialBench dataset (chen2025trialbench).
-Using 20769 annotated Phase I-IV trials, we encode multi-modal features into 6,173-dimensional embeddings (details in Appendix A.1.2) and train LightGBM (ke2017lightgbm) classifiers to predict trial outcome $\hat{y}\in[0,1]$.
-We follow TrialBench’s train-test split, further splitting the training set 8:2 for training-validation.
-Due to computational constraints, we evaluate the agent on a stratified sample of 60 failed trials from the test set (20 enrollment, 20 safety, 20 efficacy failures) representing diverse trial phases (Data detail in Appendix A.1.1).
-The agent operates with a 5-iteration budget.
-We employ an adaptive exploration strategy: when the estimated factorial space <1,000<1,000 combinations, we perform exhaustive exploration to find the optimal solution; otherwise use beam search with width $k=8$.
-We measure effectiveness through predicted probability improvement, threshold achievement rate, and convergence efficiency.
-4.2 Simulation Environment Performance
-Our Simulation Environment’s model (GBDT) is compared against Baseline approaches (TrialBench (chen2025trialbench) and HINT (fu2022hint), prior SOTA systems operating on original TrialBench features), and Logistic Regression also trained on the same encodes (Appendix A.1.2).
-Failure-specific Prediction.
-Implemented in ClinicalReTrialAgent, the simulation environment must correctly predict specific failure outcomes (enrollment, safety, efficacy).
-We train three independent binary GBDT classifiers on our encoded features, each targeting one failure detection task against success.
-Table 1, 2 and 3 report comprehensive metrics across our models and baseline approaches re-trained for the binary task: Poor Enrollment, Safety/Adverse Effect and Lack of Efficacy prediction.
-Our model achieves PR-AUC >0.75>0.75 across all failure modes, meeting the threshold for reliable discriminative feedback.
-All models achieve Failure Detection Rates of 70-74% at task-specific thresholds ($p\geq 0.6$ for enrollment, $p\geq 0.9$ for safety, $p\geq 0.85$ for efficacy).
-This ensure that predicted probability shifts $\Delta p>0.03$ reliably indicate improved trial designs (Appendix A.1.1).
-Table 1: Performance on Poor Enrollment prediction.
-Model ROC-AUC PR-AUC Fail Det.
-TrialBench 0.613 $\pm$ 0.007 0.626 $\pm$ 0.011 0.525 $\pm$ 0.013
-HINT 0.534 $\pm$ 0.010 0.613 $\pm$ 0.012 0.580 $\pm$ 0.018
-Logistic Reg. 0.622 $\pm$ 0.010 0.696 $\pm$ 0.012 0.669 $\pm$ 0.012
-GBDT (ours) 0.676 $\pm$ 0.009 0.754 $\pm$ 0.010 0.740 $\pm$ 0.012
-Table 2: Performance on Drug Adverse Effect prediction.
-Model ROC-AUC PR-AUC Fail Det.
-TrialBench 0.587 $\pm$ 0.017 0.892 $\pm$ 0.006 0.427 $\pm$ 0.035
-HINT 0.513 $\pm$ 0.014 0.882 $\pm$ 0.009 0.459 $\pm$ 0.031
-Logistic Reg. 0.612 $\pm$ 0.018 0.909 $\pm$ 0.008 0.422 $\pm$ 0.029
-GBDT (ours) 0.656 $\pm$ 0.018 0.925 $\pm$ 0.007 0.695 $\pm$ 0.028
-Table 3: Performance on Drug Efficacy prediction.
-Model ROC-AUC PR-AUC Fail Det.
-TrialBench 0.692 $\pm$ 0.012 0.862 $\pm$ 0.006 0.565 $\pm$ 0.020
-HINT 0.559 $\pm$ 0.013 0.841 $\pm$ 0.008 0.525 $\pm$ 0.021
-Logistic Reg. 0.665 $\pm$ 0.015 0.886 $\pm$ 0.009 0.549 $\pm$ 0.025
-GBDT (ours) 0.746 $\pm$ 0.013 0.914 $\pm$ 0.007 0.725 $\pm$ 0.021
-TrialBench Benchmark.
-We further validate the same model architecture against existing benchmarks on the TrialBench 4-class classification task (predicting Success, Enrollment Failure, Safety Failure, or Efficacy Failure).
-Table 4 shows the Simulation Environment’s base model in our simulation environment performed best over all baselines, with higher ROC-AUC of 0.06 to 0.19.
-Category Model Phase 1 Phase 2 Phase 3 Phase 4
-ROC-AUC PR-AUC ROC-AUC PR-AUC ROC-AUC PR-AUC ROC-AUC PR-AUC
-Baseline TrialBench 0.475±0.027 0.255±0.006 0.569±0.010 0.295±0.008 0.550±0.012 0.279±0.008 0.477±0.022 0.256±0.007
-HINT 0.540±0.022 0.272±0.009 0.535±0.009 0.267±0.005 0.474±0.019 0.251±0.006 0.548±0.021 0.273±0.014
-ClinicalReTrial Logistic Reg. 0.606±0.019 0.326±0.015 0.583±0.011 0.306±0.008 0.621±0.016 0.350±0.017 0.550±0.022 0.280±0.010
-GBDT 0.633±0.016 0.344±0.016 0.662±0.011 0.382±0.011 0.669±0.017 0.412±0.019 0.543±0.025 0.282±0.012
-Table 4: Performance of multi-class clinical trial outcome prediction across trial phases.
-Feature Importance Analysis.
-We studied feature importance analysis with SHAP (lundberg2017shap).
-Consistent with our hypothesis, sentence-level eligibility, Drug-disease interaction features and endpoint alignment features are most important for outcomes prediction (Appendix A.1.3).
-4.3 Agent’s Protocol Optimization
-Having confirmed the simulation environment’s reliability, we evaluate ClinicalReTrialAgent’s ability to redesign failed clinical trials.
-4.3.1 Convergence Analysis
-Table 5 reports comprehensive convergence statistics across all trials.
-The Agent had 83.3% of protocol designs improved (50/60 successfully processed trials showed positive $\Delta p$), with 4 trials (6.7%) encountering agent failures where the system identified zero opportunities of potential redesign, occurring in the efficacy failure mode.
-Table 5: Convergence analysis across 60 test trials.
-Mode Trials Failures Pos. $\Delta p$ Threshold
-Enrollment 20 0 20/20 (100%) 8/20 (40%)
-Safety 20 0 18/20 (90%) 4/20 (20%)
-Efficacy 20 4 12/20 (60%) 10/20 (50%)
-Overall 60 4 (6.7%) 50/60 (83.3%) 22/60 (36.7%)
-The system demonstrated efficient convergence patterns, with 15% (9/60) of trials exhibiting natural termination before iteration 5 due to exhausted modification space.
-Most trials used all 5 iterations, suggesting adaptive stopping could improve efficiency.
-4.3.2 Trial Redesign Learning Dynamics
-We examine the learning trajectory of successfully processed trials (56/60).
-The system achieved mean improvement of $\Delta p=+0.057$ (95% CI: [+0.040, +0.074]).
-Table 6 stratifies results by failure mode, while Figure 2 illustrates learning dynamics by iteration.
-Table 6: Probability shift ($\Delta p$) analysis by failure mode across 56 trials. IQR is 25th-75th percentile range.
-Mode Trials $p_{0}$ / $p_{final}$ $\Delta p$ IQR
-Enrollment 20 0.506 / 0.563 +0.058 [+0.034, +0.068]
-Safety 20 0.791 / 0.831 +0.070 [+0.032, +0.092]
-Efficacy 16 0.813 / 0.859 +0.040 [+0.013, +0.039]
-Overall 56 0.695 / 0.730 +0.057 [+0.029, +0.073]
-Figure 2: Iterative learning dynamics stratified by failure mode. IQR bars span the 25th–75th percentiles (interquartile range) across runs.
-Performance heterogeneity across failure modes reflects the differential amenability of clinical trial design elements to protocol-level intervention.
-Safety failures exhibit the largest improvements (mean $\Delta p=+0.070$, IQR [+0.032, +0.092]), as adverse events often stem from identifiable contraindication patterns that can be systematically addressed through eligibility refinement and dosage adjustment.
-Enrollment failures show substantial gains (mean $\Delta p=+0.058$), consistent with the observation that recruitment barriers frequently arise from overly restrictive or poorly specified inclusion criteria rather than fundamental feasibility constraints.
-Efficacy failures demonstrate the smallest yet statistically meaningful improvements (mean $\Delta p=+0.040$), as therapeutic effectiveness depends heavily on drug-disease compatibility, where sometimes protocol modifications alone cannot fix the essential drug failure.
-As shown in Figure 2, the learning trajectory reveals major initial gains followed by decreasing returns.
-This pattern validates the knowledge distillation mechanism: high-quality modifications are identified early through rapid retrieval augmented analysis, while later iterations exploit narrower optimization opportunities by refining secondary parameters or addressing edge case contraindications.
-4.3.3 Computational Efficiency
+9:  $\mathcal{M}_{t}^{\text{local}}\leftarrow\mathcal{M}_{t-1}^{\text{local}}\cup\{\mathcal{K}_{t},\mathcal{R}_{t},\mathcal{H}_{t}\}$
+10:end for
 […]
-Table 7: Computational efficiency by failure mode.
+12:return $T^{*}$, $r_{\text{best}}$
+Appendix B Computation Cost
+Table 5: Computational cost by failure mode.
 […]
@@
 Overall 56 0.122 +0.053 3.68
-4.4 Ablation Study on Self-improvement
-To validate architectural components, we conducted paired ablation across 10 enrollment failure trials, systematically removing: (1) memory-guided iterative learning and (2) redesign pool optimization.
-Each trial was evaluated under all three conditions with identical initialization, enabling within-subjects comparison.
-Both components contribute significantly and independently (Figure 3, Table 8).
-Memory removal degraded performance at iteration 1 ($\Delta p=+0.0131$, $p=0.042$) and iteration 5 ($\Delta p=+0.0190$), demonstrating immediate warm start benefits that strengthen over iterations.
-Removing the redesign pool yielded comparable degradation ($\Delta p=+0.0126$).
-These findings validate our design: memory enables the agent to manage intelligent exploration, while the redesign pool enables exploitation by reusing successful modifications.
-Figure 3: Self-Improving ablation study across 10 trials.
-Full system (blue) outperforms setups without memory (red) or redesign pool optimization (green).
-Memory provides early benefits, and pool effects compound over iterations.
-IQR bars span the 25th–75th percentiles across runs.
-Figure 4: ClinicalReTrial Agent’s flowchart on Poor Enrollment failed trial case study (NCT01298752, 2011-02-16), together with the real-world redesign (NCT01591161, 2012-05-02).
-The Agent’s iterative refinement of failure analysis and according modifications are demonstrated.
-Table 8: Self-Improving ablation study. Stats computed using paired $t$-tests with $n=10$ trials. *$p$-value <0.05<0.05, **$p$-value <0.01<0.01.
-Removed Iteration Ablation Full System $\Delta p$ $p$-value Cohen’s $d_{z}$
-Memory 1 $024\pm 0.016 0.0$$37\pm 0.020 +0.0$13+0.013 $0.042^{*}$ $0.81$
-Memory 5 $038\pm 0.016 0.0$$57\pm 0.023 +0.0$19+0.019 $0.007^{**}$ $1.10$
-Pool 5 $045\pm 0.022 0.0$$57\pm 0.023 +0.0$13+0.013 $0.005^{**}$ $1.18$
-4.5 Retrospective Case Studies: Real-World Validation
-To validate ClinicalReTrial’s clinical applicability, we analyze trial pairs, where investigators successfully redesigned and re-executed failed protocols spanning enrollment, safety, and efficacy failure modes (Appendix A.5).
-Validating agent reasoning against real-world protocol redesigns provides critical insight into clinical applicability.
-Here we present a poor enrollment redesign case (safety and efficacy cases in Appendix A.5): NCT01298752, a Phase 3 trial of Mapracorat (anti-inflammatory ophthalmic suspension) for post-cataract surgery inflammation that failed due to slow enrollment.
-Sponsored by Bausch & Lomb, the trial was subsequently redesigned and successfully executed as NCT01591161.
-Figure 4 illustrates ClinicalReTrial’s iterative refinement process across three optimization cycles.
-The Enrollment barrier (cataract surgery waiting requirement) is efficiently identified with positive reward provided by simulation environment.
-The agent also progressively explores the modification space: baseline AC cell requirements is successfully added as an enrichment criterion; while agent also explores the safety enhancement, but end up failing to align with real-world redesign.
-5 Conclusion
-In this work, we have presented ClinicalReTrial, a novel self-evolving agent framework that moves beyond passive clinical trial outcome prediction to enable proactive optimization of clinical trial protocols.
-By integrating an interpretable, accurate simulator with an autonomous Agent capable of causal reasoning and iterative design refinement, our system not only able to forecast trial success but also generates actionable modifications, including adjusting eligibility criteria, dosing regimens, or endpoint definitions to enhance feasibility and likelihood of success.
-Evaluated on standardized TrialBench benchmark, ClinicalReTrial achieves strong predictive performance while demonstrating the ability to discover significant protocol improvements with clinical best practices.
-6 Limitations
-Our framework has several limitations that suggest directions for future research.
-First, the simulation environment’s predictive accuracy creates potential for false improvement signals and missed opportunities, though validation filtering and iterative refinement provide partial mitigation; this constraint could be addressed by integrating future state-of-the-art prediction models as drop-in replacements for the GBDT oracle.
-Second, the system lacks adaptive convergence detection; 83.9% of trials exhausted the full 5-iteration budget rather than stopping when modification potential plateaus, suggesting the need for learned stopping criteria based on diminishing returns patterns.
-Third, retrospective case study analysis reveals tactical domain knowledge gaps: while the system excels at strategic-level reasoning, it may struggle with operational specifics such as selecting appropriate biomarkers, anticipating implementation constraints, and distinguishing when radical simplification outperforms incremental fortification, often over-relying on complexity multiplication where parsimony proves more effective.
-Future work should prioritize prospective validation in collaboration with clinical trial sponsors, integration of specialized biomarker knowledge bases to address tactical gaps, development of adaptive stopping mechanisms to improve computational efficiency, and expansion to larger-scale evaluation encompassing broader disease areas and trial designs.
-Appendix A Appendix
-A.1 Simulation Environment Details
-A.1.1 Dataset Statistics
-Dataset we used to train the prediction models comprises 20,769 clinical trials from TrialBench’s failure reason dataset. Table 9 shows the label distribution across four categories.
+Appendix C Simulation Environment Details
+C.1 Dataset Statistics
+Dataset used for training the prediction models comprises 20,769 clinical trials from TrialBench’s failure reason dataset. Table 6 shows the label distribution across four categories.
 Table 6: Distribution of failure reason labels in the dataset.
@@
 Total 60 100
-A.1.2 Encode Details
+C.2 Encode Details
 This appendix provides comprehensive implementation details for the Simulation Environment described in §3.2.4, including encoder pretraining procedures, model training hyperparameters, and detailed validation results.
 Text Features.
-Textual contents are encoded using BioBERT (lee2020biobert), a domain-adapted language model pre-trained on PubMed abstracts and PMC full-text articles.
+Textual contents are encoded using BioBERT (Lee et al., 2020), a domain-adapted language model pre-trained on PubMed abstracts and PMC full-text articles.
 Critically, we diverge from prior work by decomposing eligibility criteria at the sentence level rather than treating them as monolithic text blocks.
@@
 $\mathbf{e}_{s_{i}}=\text{BioBERT}(s_{i}),\quad\mathbf{h}_{\mathcal{T}}=\max_{i=1}^{n}\mathbf{e}_{s_{i}}$ (2)
-This sentence-level representation preserves granularity essential for aspect-specific modification—ClinicalReTrialAgent can target individual criteria rather than generic protocol summaries.
+This sentence-level representation preserves granularity essential for aspect-specific modification: ClinicalReTrial Agent can target individual criteria rather than generic protocol summaries.
 Graph Features.
@@
 $\mathbf{h}_{\text{drug}}=\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}\mathbf{h}_{u}\in\mathbb{R}^{d_{\text{mpnn}}}$ (5)
-For trials with multiple drugs, we average their embeddings. The MPNN encoder is pretrained on pharmacokinetic (ADMET) tasks, then fine-tuned on trial outcome labels (details in Appendix A.1).
+For trials with multiple drugs, we average their embeddings. The MPNN encoder is pretrained on pharmacokinetic (ADMET) tasks, then fine-tuned on trial outcome labels (details in Appendix C).
 Disease Hierarchical Encoding.
@@
 For trials targeting multiple diseases, we average their embeddings.
-The GRAM encoder is initialized with the ICD-10 hierarchical ontology, then fine-tuned on historical trial success rates (details in Appendix A.1).
+The GRAM encoder is initialized with the ICD-10 hierarchical ontology, then fine-tuned on historical trial success rates (details in Appendix C).
 Tabular Features.
@@
 We train models separately for each failure mode, enabling ClinicalReTrial Agent to target specific causes during protocol optimization.
-Our experiments utilize the TrialBench dataset (chen2025trialbench), which contains over 12,000 annotated clinical trials spanning Phase I through Phase IV, with each trial labeled according to outcome.
+Our experiments utilize the TrialBench dataset (Chen et al., 2025), which contains over 12,000 annotated clinical trials spanning Phase I through Phase IV, with each trial labeled according to outcome.
 The dataset provides multi-modal features including drug molecular structures, disease ICD-10 codes, eligibility criteria text, trial metadata, and intervention details.
@@
 where semicolons denote concatenation.
-For each failure mode $\tau\in\{\text{enrollment, safety, efficacy}\}$, we train a separate LightGBM classifier $\mathcal{M}_{\tau}$ that predicts trial success probability, optimizing binary cross-entropy loss.
+For each failure mode $\tau\in\{\text{enrollment, safety, efficacy}\}$, we train a separate LightGBM classifier $\mathcal{M}_{\tau}$ that predicts trial success probability.
 The predicted probability $\hat{y}=\mathcal{M}_{\tau}(\mathbf{x}_{\text{trial}})\in[0,1]$ serves as the reward signal for evaluating protocol modifications in the agent system.
 Model Training and Validation.
-We employ LightGBM (ke2017lightgbm) for its computational efficiency with high-dimensional sparse features.
+We employ LightGBM (Ke et al., 2017) for its computational efficiency with high-dimensional sparse features.
 Three independent models are trained for enrollment, safety, and efficacy failure prediction using cross-validation with early stopping.
 The trained GBDT models achieve strong predictive performance across all failure modes (PR-AUC >> 0.75) with well-calibrated probability estimates, validating the simulation environment as a reliable proxy for real trial outcomes.
-A.1.3 Ablation Study
+C.3 Ablation Study
 Word-Level Attention Analysis.
@@
 Figure 6: Feature importance of each encoder on 3 classification tasks (enrollment, safety, and efficacy), measured by PR-AUC drop when the encoder is masked out during prediction.
-A.2 Analysis Agent Details
-The Analysis Agent implements a domain-aware ReAct reasoning pipeline adapted by failure mode (enrollment, safety, efficacy).
+Appendix D Validator Case Studies
+Case 1: Disease-RAG prevents misaligned safety filtering in a Type 2 diabetes trialcolback Clinical trial case.
+Trial: NCT00294723.
+Condition: Type 2 diabetes (liraglutide safety evaluation).
+Proposed redesign: exclude patients with eGFR <60<60 mL/min at screening.
+Retrieved evidence (Disease-RAG).
+Query: Diabetes.
+Retrieved guidelines emphasized management of cardiovascular disease risk factors as a core part of diabetes care.
+Validation outcome.
+Decision: Invalid redesign.
+Reason.
+Combined with the adverse event summary, which included myocardial infarctions, the retrieved disease context showed that the proposed renal exclusion was overly broad and not well aligned with the trial’s main safety risks.
+The Validator Agent therefore pruned the modification.
+Case 2: PubMed-RAG prevents unsafe diagnostic relaxation in a hypertension-in-pregnancy trialcolback Clinical trial case.
+Trial: NCT03872336.
+Condition: Hypertension in pregnancy (labetalol treatment).
+Proposed redesign: accept a single severe blood pressure reading ($\geq 160/110$ mmHg) for diagnosis.
+Retrieved evidence (PubMed-RAG).
+Query: Hypertension diagnosis measurement criteria.
+PMID 29574470 and PMID 24652215 both support diagnosis based on repeated blood pressure measurements over time.
+Validation outcome.
+Decision: Invalid redesign.
+Reason.
+The Validator Agent determined that accepting a single severe reading is inconsistent with standard diagnostic practice and may cause false-positive diagnoses, unnecessary intervention, or inappropriate enrollment.
+Case 3: Drug-RAG preserves a critical safety exclusion in a labetalol trialcolback Clinical trial case.
+Trial: NCT03872336.
+Condition: Hypertension in pregnancy (labetalol treatment).
+Proposed redesign context: remove or weaken the exclusion criterion on persistent mild–moderate asthma.
+Retrieved evidence (Drug-RAG).
+Query: labetalol toxicity.
+Retrieved safety evidence noted risk of hypotension, bradycardia, and bronchospasm.
+Validation outcome.
+Decision: Preserve existing criterion.
+Reason.
+The Validator Agent identified a drug–disease interaction between labetalol and asthma risk.
+Because bronchospasm is a foreseeable adverse effect, removing the asthma exclusion would be clinically unsafe.
+Appendix E Analyzer Agent Details
+The Analyzor Agent implements a domain-aware ReAct reasoning pipeline adapted by failure mode (enrollment, safety, efficacy).
 Novel components include adverse event profiling, statistical power assessment, and design-level pivots.
 Table 10 summarizes failure-mode-specific adaptations.
-Table 13: Analysis Agent variants by failure mode.
+Table 10: Analyzer Agent variants by failure mode.
 Component Enrollment Safety Efficacy
@@
 </dosage_tradeoff>
-A.3 Augmentation Agent Details
-Table 14: Augmentation Agent novel features by failure mode. Process differ a little in dosage modification strategy across failure modes.
+Appendix F Generator Agent Details
+Table 11: Generator Agent novel features by failure mode. Process differ a little in dosage modification strategy across failure modes.
 Feature Enrollment Safety Efficacy
@@
 Domain Focus Enrich participation Tighten safety exclusions Add biomarker enrichment
-A.3.1 Few-Shot Learning Mechanism
+F.1 Few-Shot Learning Mechanism
 Few-Shot Example Injection (Shared Structure) Matching Logic:
@@
 Effect: LLM learns from previous iteration’s successes/failures. Only available iteration 2+ after prev_rules established.
-A.3.2 Augmentation Prompts by Failure Mode
+F.2 Modification Prompts by Failure Mode
 Eligibility Example Role: Clinical researcher generating criterion variations Task: Generate num_augment variations with few-shot guidance Input:
@@
 </augmentations>
-A.4 Agent Output Template
+Appendix G Agent Output Template
 This section presents the structured output format produced by the agent pipeline.
@@
           "[Modified value 3]"
-A.5 Case Study Details
+Appendix H Case Study Details
 We validate ClinicalReTrial Agent’s reasoning against real-world protocol modifications provides critical insight into clinical applicability.
@@
 This suggests that future work should integrate specialized biomarker databases and safety constraint ontologies to bridge the gap between strategic reasoning and actionable clinical knowledge.
+Appendix I Ethics Statement
+I.1 Potential Risks
+This work focuses on developing AI systems to optimize clinical trial protocols through simulation-based evaluation.
+While the system demonstrates potential to improve trial design efficiency, we acknowledge several important limitations and risks:
+Decision Support, Not Replacement: ClinicalReTrial is designed as a decision support tool for clinical trial designers and should not replace human expert judgment.
+All system-generated protocol modifications require review by qualified medical professionals and regulatory compliance verification before real-world implementation.
+Simulation Environment Limitations: Our prediction models achieve PR-AUC > 0.75, but prediction errors could lead to suboptimal redesign recommendations.
+The system’s suggestions should be validated through standard clinical trial design processes and regulatory review.
+Retrospective Validation: Our case studies demonstrate alignment with real-world redesigns but are retrospective analyses.
+Prospective validation in collaboration with clinical trial sponsors is necessary before deployment.
+Generalization Constraints: The system is trained on historical clinical trial data and may not generalize to novel therapeutic mechanisms, rare diseases, or emerging trial paradigms not well-represented in the training data.
+I.2 Data Consent
+This study exclusively utilizes publicly available datasets that do not require additional consent:
+  • TrialBench Dataset (Chen et al., 2025): Publicly released benchmark containing anonymized clinical trial protocols from ClinicalTrials.gov
+  • ClinicalTrials.gov: Public registry of clinical trials maintained by the U.S. National Library of Medicine
+  • PubMed: Public database of biomedical literature abstracts
+  • DrugBank (Wishart et al., 2018): Publicly available bioinformatics and cheminformatics database
+  • Disease Database (Chen et al., 2024a): Publicly available disease ontology database
+All data sources are publicly accessible and designed for research purposes.
+No patient-level identifiable information is used in this study.
+Clinical trial protocols contain only de-identified, aggregate information as required by ClinicalTrials.gov data sharing policies.
+I.3 Ethics Review Board Approval
+This computational study analyzes publicly available, de-identified clinical trial metadata and does not involve human subjects research, prospective clinical interventions, or collection of new patient data.
+The retrospective case studies (§4.4) analyze publicly registered clinical trials with outcomes already recorded in ClinicalTrials.gov, constituting secondary analysis of publicly available data exempt from human subjects research requirements.
+Appendix J Use of Large Language Model
+Within our data construction workflow, we utilize large language models for agent in-context learning, reasoning, and augmentations generating.
+Additionally, we employ LLMs such as ChatGPT to help improve the clarity and fluency of our written content.